### v1

# Global Convergence of Policy Gradient for Entropy Regularized Linear-Quadratic Control with multiplicative noise

Abstract
Reinforcement Learning (RL) has emerged as a powerful framework for sequential decision-making in dynamic environments, particularly when system parameters are unknown. This paper investigates RL-based control for entropy-regularized Linear Quadratic control (LQC) problems with multiplicative noises over an infinite time horizon. First, we adapt the Regularized Policy Gradient (RPG) algorithm to stochastic optimal control settings, proving that despite the non-convexity of the problem, RPG converges globally under conditions of gradient domination and near-smoothness. Second, based on zero-order optimization approach, we introduce a novel model free RL algorithm: Sample-Based Regularized Policy Gradient (SB-RPG). SB-RPG operates without knowledge of system parameters yet still retains strong theoretical guarantees of global convergence. Our model leverages entropy regularization to accelerate convergence and address the exploration versus exploitation trade-off inherent in RL. Numerical simulations validate the theoretical results and demonstrate the efficacy of SB-RPG in unknown-parameters environments.
1 Introduction
Reinforcement Learning (RL) is a subfield of machine learning that focuses on training agents to make sequential decisions by interacting with dynamic environments. Unlike supervised learning, which relies on labeled datasets, RL agents learn through trial and error, guided by a reward signal that quantifies the desirability of their actions [1]. The ultimate goal is to discover an optimal policy—a mapping from states to actions which maximizes the cumulative long-term rewards. In recent years, RL has revolutionized the field, achieving human-level performance in domains ranging from game playing [2] to robotics [3], and autonomous driving [4].
Optimal control theory seeks to design control policies that maximize a predefined performance criterion for dynamic systems. RL and optimal control are naturally aligned in their fundamental principles as both approaches incorporate decision-making considerations. However, traditional optimal control requires complete knowledge of all environmental parameters to make decisions. Obtaining such precise parameters is infeasible in real-world scenarios, making the application of optimal control particularly challenging. RL based control in situations where the system parameters are unknown has achieved significant success in recent years. Linear-quadratic (LQ) control problem, as one of the most fundamental problems in control theory, has attracted considerable attention and has been extensively studied in the RL based control literature. For example, in the continues time setting, Wang and Zhou [5] adopt the RL method to solve mean–variance portfolio problem, Wang et,al. [6] carry out a complete theoretical analysis of RL based LQ control problem, Li et al. [7] employ a policy iteration RL approach to investigate LQ mean-field control problems over an infinite horizon.
Policy Gradient [8] is a class of RL algorithm. It directly parameterizes the optimal policy and performs gradient descent on the policy, which makes it easy to implement and widely applicable. However, even for the most basic LQ control problem, policy gradient methods encounter a non-convex optimization landscape [9]. As a result, the convergence properties of policy gradient algorithms is a significant area of research, and numerous useful results are obtained in this field. Fazel et al. [9] prove the global convergence in the time homogeneous, infinite time horizon deterministic LQ problem. Based on [9], Gravell et al. [10] [11] extend the global convergence results to LQ with multiplicative noise through rigorous proofs, and Lai et al. [12] extend it to LQ with both multiplicative and additive noises. In the finite time setting, Hambly et al. [13] consider the convergence of policy gradient methods of LQ with additive noises and show a financial application. All of the aforementioned papers are generally composed of two parts. The first part consider global convergence of policy gradient under the assumption that the system parameters are known, which corresponds to the model-based RL (i.e., estimating the model parameters and learning). While the second part builds upon the first to analyze convergence under unknown parameters setting, which corresponds to the model-free RL (i.e., end-to-end). Hu et al. [14] survey several recent theoretical advances regarding the optimization landscape, global convergence properties, and sample complexity of gradient-based methods applied to different control problems, including but not limited to linear-quadratic (LQ) systems.
A important topic in RL is the exploration-exploitation trade-off: the agent must balance exploiting known information to maximize immediate reward and exploring the environment by trying random actions in order to find potentially better actions and states. In most cases, exploration is highly resource-intensive, therefore, numerous solutions are proposed to address exploration-exploitation trade-off. The trade off between exploration and exploitation has been thoroughly studied for the LQR. For example, in [15] and is improved upon in [16]. In the discrete action space setting, $\epsilon$-greedy policy [17] and Botzman (Softmax) policy [18] are two effective and popular ways to balance exploration-exploitation and numerous developments have been made based on them. In addition to these, recent research has introduced entropy-regularized RL formulations. This approach explicitly integrates exploration into the optimization objective by including entropy as a regularization term, thereby imposing a trade-off weight on the entropy of the exploration strategy . Ahmed et al. [19] demonstrate that, even when the exact gradient is available, policy optimization remains challenging because of the complex geometry of the objective function. Moreover, employing policies with higher entropy can smooth the optimization landscape, facilitating connections between local optima. Neu et, al. [20] propose a general framework for entropy-regularized average-reward RL in Markov decision processes. It is noteworthy that while entropy-regularization has been quite useful it comes with the caveat that there are many other ways to promote policy exploration nor is entropy-regularization always effective as seen in [21].
In the context of entropy-regularized RL formulation for LQ problems, there has also been substantial research progress; however, most of these results pertain to actor-critic methods, such as the previously mentioned [5] [6] [7]. In contrast, studies focusing on policy gradient based approaches for entropy-regularized LQ problems remain relatively limited. Michael et, al. [22] study the global linear convergence of policy gradient methods for finite-horizon continuous-time entropy-regularized LQ control problems. Guo et, al. In the discrete time setting, Guo et, al. [23] propose and analyzes two new policy gradient based RL method for entropy-regularized LQ problems: regularized policy gradient (RPG) and iterative policy optimization (IPO) and prove their fast convergence given exact model parameters(i.e., model based). However, they only consider additive noise and conducted their analysis solely in the model-based setting, which limits the applicability of their methods. Multiplicative noise models may be produce more robust policies, they are still more complex than the typical additive model which may result in slower convergence. To the best of our knowledge, addressing the case of multiplicative noise in both model-based and model-free settings still remains an open problem.
Our contribution
This paper makes two fundamental contributions: First, We extend Regularized Policy Gradient (RPG) [23] method to stochastic optimal control scenarios and show that while the stochastic LQ is non-convex, the RPG still converges to a global minimum due to the property of gradient domination and almost smoothness, thereby enhancing the system’s robustness and significantly broadening its potential applications. Secondly, and more importantly, we employ a zero order optimization technique; we propose Sample Based Stochastic Regularized Policy Gradient (SB-RPG) method and rigorously prove its global convergence properties. SB-RPG does not require knowledge of the specific system parameters values. This theoretical guarantee enables our model to operate effectively in parameter-unknown scenarios, where as RPG cannot. Numerical simulations also support the theoretical results.
Notations
We adopt standard mathematical notation throughout this paper. For any matrix $Z\in\mathbb{R}^{n\times m}$, we denote $\|Z\|$ as the spectral norm of $Z$, $\|Z\|_{F}$ as the Frobenius norm of $Z$, $\sigma_{\min}(Z)$ and $\sigma_{\max}(Z)$ as the minimum and maximum singular values of $Z$, respectively, and $Z^{\top}$ as the transpose of $Z$.
Organization
For the sake of brevity, only the proofs of the main theorems are presented in the main text, while all detailed proofs of the lemmas are relegated to the Appendix. The rest of sections are organized as follows. In section 2 we formulate the optimal control problem and transform it into a optimization problem. It is natural to consider first order method to cope with optimization problem so we give the explicit form of gradient with respect to optimization variables. In section 3, we consider Regularized Policy Gradient method proposed in [23] and provide the guarantee of global convergence. In section 4, we consider the case where all the parameters are unknown, we proposed Sample Based Regularized Policy gradient (SB-RPG) to cope with this situation. In section 5, we provide numerical experiments showing the effectiveness of our algorithm.
2 Formulation
In this section, we clearly formulate the stochastic optimal control problem over an infinite time horizon with a constant discounted rate $\gamma$ and derived the optimal feedback control policy in (5). Inspired by the structure of feedback control policy, we linearly parameterize our policy. By doing so, we transformed the optimal control problem into an finite dimensional (i.e., $n$-dimensional) optimization problem. Since optimization naturally involves consideration of first-order derivatives, we provide the explicit form of the first-order derivative.
Consider the following discrete time exploratory stochastic linear quadratic control system in the infinite time horizon:
$x_{t+1}$ $=(A+w_{t}^{x}C)x_{t}+(B+w_{t}^{u}D)u_{t}$ (1)
where $A,C,x_{t},w_{t}^{x}\in\mathbb{R},~B,w_{t}^{u}\in\mathbb{R}^{1\times n},~D\in\mathbb{R}^{n\times n},$ and $w_{t}^{x},w_{t}^{u}$ are white noises, which are distributed as follows
$\mathbb{E}[w^{x}_{t}]=0,~~\mathbb{E}[(w_{t}^{x})^{2}]=1$
$\mathbb{E}[w^{u}_{t}]=\mathbf{0}_{1\times n},~~\mathbb{E}[(w_{t}^{u})^{\top}w_{t}^{u}]=I_{n\times n}$
Unlike traditional optimal control problems that focus solely on deterministic control, we incorporate entropy regularization and consider stationary randomized Markovian policies. This approach enables us to effectively address the exploration problem in Reinforcement Learning. Specifically, we define the set of admissible policies as $\Pi:=\{\pi:\mathcal{X}\to\mathcal{P}(\mathcal{U})\}$, where $\mathcal{X}$ denotes the state space, $\mathcal{U}$the action space, and $\mathcal{P}(\mathcal{U})$ the set of probability measures over $\mathcal{U}$. Each admissible policy $\pi\in\Pi$ assigns to every state $x\in\mathcal{X}$ a probability distribution over actions in $\mathcal{U}$.
For any given policy $\pi\in\Pi$, the associated Shannon entropy is defined as
$\mathcal{H}(\pi(\cdot|x)):=-\int_{\mathcal{U}}\pi(u|x)\log\pi(u|x)\,du,$
which measures the uncertainty or information gain from exploring the environment. By incorporating this entropy term as a regularization component in the objective function, we encourage the policy to gather information about the unknown environment and to promote exploration. The objective functional then takes the following form:
$\min_{\pi\in\Pi}\,\mathbb{E}_{x\sim\mathcal{D}}[J(x)],$ (2)
where $\Pi$ is the admissible policy set and
$J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=0}^{\infty}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{T}Ru_{t}-\tau\mathcal{H}(\pi(\cdot|x_{t}))\right)\right|x_{0}=x\right]$ (3)
Now that we have defined the exploratory stochastic LQ problem, we present the following theorem, which provides the optimal policy, the optimal objective value, and the corresponding Algebraic Riccati Equation (ARE).
Theorem 2.1 (Optimal value functions and optimal control).
The optimal value function $J^{*}:\mathcal{X}\rightarrow\mathbb{R}$ in can be expressed as $J^{*}(x)=Px^{2}+q$ with $P$ satisfying the following Algebraic Riccati Equation (ARE)
$P$ $=Q+\gamma P(A^{2}+C^{2})-(\gamma AP)^{2}B(R+\gamma P(B^{\top}B+D^{\top}D))^{-1}B^{\top},$
$q$ $=\frac{\text{Tr}(\Sigma^{*}R)+\gamma P\text{Tr}(\Sigma^{*}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(k+\log((2\pi)^{k}\operatorname{det}\Sigma^{*})}{1-\gamma},$
where
$K^{*}=\gamma(R+\gamma P(B^{\top}B+D^{\top}D))^{-1}APB^{\top},\quad\Sigma^{*}=\frac{\tau}{2}(R+\gamma P(B^{\top}B+D^{\top}D))^{-1},$ (4)
for any $x\in\mathcal{X}$, the corresponding optimal policy for system (1) and objective functional (2) is :
$\pi^{*}=\mathcal{N}(-K^{*}x,\Sigma^{*}).$ (5)
The proof of Theorem 2.1 relies on the following lemma, which establishes the optimal solution for the one-step reward function in the presence of entropy regularization. This lemma provide the necessary foundation for deriving both the optimal policy and the corresponding value function under entropy-regularized rewards. The proof of Lemma 2.1 is provided in Section 8.1 of [23].
Lemma 2.1.
For any given symmetric positive definite matrix $M\in\mathbb{R}^{k\times k}$ and vector $b\in\mathbb{R}^{k}$, the optimal solution $p^{*}\in\mathcal{P}(\mathcal{U})$ to the following optimization problem is a multivariate Gaussian distribution with covariance $\frac{\tau}{2}M^{-1}$ and mean $\frac{1}{2}M^{-1}b$:
$\min_{p\in\mathcal{P}(\mathcal{U})}$ $\mathbb{E}_{u\sim p(\cdot)}\left[u^{T}Mu+b^{T}u+\tau\log p(u)\right],$
subject to $\int_{\mathcal{U}}p(u)du=1,$
$p(u)\geq 0,\quad\forall u\in\mathcal{U}.$
Proof.
(of Theorem 2.1). By definition of $J^{*}$ in (2),
$J^{*}(x)=\min_{\pi\in\Pi}\mathbb{E}_{\pi}\Bigl\{Qx^{2}+u^{T}Ru+\tau\log(\pi(u|x))+\gamma J^{*}((A+w_{t}^{x}C)x_{t}+(B+w_{t}^{u}D)u_{t})\Bigr\},$ (6)
where the expectation is taken with respect to $u\sim\pi(\cdot|x)$ and the noise terms $w_{t}^{u}$ and $w_{t}^{x}$, with mean 0 and covariance $I_{n\times n}$. Stipulating
$J^{*}(x)=Px^{2}+q$ (7)
for a positive $P,q\in\mathbb{R}$ and plugging into (6), we can obtain the optimal value function with dynamic programming principle:
$J^{*}(x)$ $=Qx^{2}+\min_{\pi}\mathbb{E}_{\pi}\Bigl\{u^{T}Ru+\tau\log(\pi(u|x))$
$\quad+\gamma\left[P((A+w^{x}C)x+(B+w^{u}D)u)^{2}+q\right]\Bigr\}$
$=(Q+\gamma P(A^{2}+C^{2}))x^{2}+\gamma q$
$\quad+\min_{\pi}\mathbb{E}_{\pi}\Bigl\{u^{\top}(R+\gamma P(B^{\top}B+D^{\top}D))u+\tau\log(\pi(u|x))+2\gamma APxBu\Bigr\}.$
Now apply Lemma 2.1 to (3) with $M=R+\gamma P(B^{\top}B+D^{\top}D)$ and $b=2\gamma APxB^{\top}$, we can get the optimal policy at state $x$:
$\pi^{*}(\cdot|x)$ $=\mathcal{N}\left(-(R+\gamma P(B^{\top}B+D^{\top}D))^{-1}\gamma APxB^{\top},\frac{\tau}{2}(R+\gamma P(B^{\top}B+D^{\top}D))^{-1}\right)$ (8)
$=\mathcal{N}\left(-K^{*}x,\Sigma^{*}\right),$
where $K^{*},\Sigma^{*}$ are defined in (8). To derive the associated optimal value function, we first calculate the negative entropy of policy $\pi^{*}$ at any state $x\in\mathcal{X}$:
$\mathbb{E}_{\pi^{*}}[\log(\pi^{*}(u|x))]=\int_{\mathcal{A}}\log(\pi^{*}(u|x))\pi^{*}(u|x)du=-\frac{1}{2}\left(k+\log\left((2\pi)^{k}\operatorname{det}\Sigma^{*}\right)\right).$ (9)
Plug (8) and (9) into (3) to get
$J^{*}(x)$ $=(Q+\gamma P(A^{2}+C^{2}))x^{2}+\gamma q$
$\quad+\min_{\pi}\mathbb{E}_{\pi}\Bigl\{u^{\top}(R+\gamma P(B^{\top}B+D^{\top}D))u+\tau\log(\pi(u|x))+2\gamma APxBu\Bigr\}$
$=(Q+\gamma P(A^{2}+C^{2}))x^{2}+\gamma q-\frac{\tau}{2}(k+\log((2\pi)^{k}\operatorname{det}\Sigma^{*})$
$\quad+Tr(\Sigma^{*}(R+\gamma P(B^{\top}B+D^{\top}D))$
$\quad+(K^{*}x)^{\top}\ (R+\gamma P(B^{\top}B+D^{\top}D))(K^{*}x)-{2\gamma APxBK^{*}x}$
$=x^{2}[Q+\gamma P(A^{2}+C^{2}-ABK^{*})]$
$\quad-\frac{\tau}{2}(k+\log((2\pi)^{k}\operatorname{det}\Sigma^{*})+\gamma q+Tr(\Sigma^{*}(R+\gamma P(B^{\top}B+D^{\top}D))$
Combining this with (7), the proof is completed. ∎
Inspired by the form of the optimal control function, we can linearly parameterize our policy and transform the optimal control problem to the optimization problem by the parameters $(K,\Sigma)$. By doing so, the policy can be formulated as $\pi(u|x)=\mathcal{N}(-Kx,\Sigma)$. Then admissible policy set for $(K,\Sigma)$ is defined as $\Omega=\{K\in\mathbb{R}^{n},\Sigma\in\mathbb{R}^{n\times n}:\gamma V_{K}<1,\Sigma\succ 0,\Sigma^{\top}=\Sigma\}$. For simplicity of notation, we define the cost of system given the deterministic initial state $x_{0}$ as $C_{K,\Sigma}(x_{0}){:=}\mathbb{E}\left.\big[\sum_{t=0}^{\infty}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{T}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)\right|u_{t}\sim\mathcal{N}(-Kx_{t},\Sigma)\big]$.
Lemma 2.2 (Optimization formulation).
The optimal control problem consider in Theorem 2.1 can be written as follows:
$\min_{(K,\Sigma)\in\Omega}f(K,\Sigma)=\mathbb{E}_{x_{0}\sim\mathcal{D}}[C_{K,\Sigma}(x_{0})]=P_{K}\mu+q_{K,\Sigma}$ (10)
where $\pi(u|x)=\mathcal{N}\left(-Kx,\Sigma\right)$, $\mu=\mathbb{E}_{x_{0}\sim\mathcal{D}}x^{2}_{0}$ and $x_{t}$ subject to the dynamics of system in (1) and $P_{K},q_{K,\Sigma}$ satisfy the following functions:
$P_{K}=Q+K^{\top}RK+\gamma P_{K}(A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK)$ (11)
$q_{K,\Sigma}=\frac{\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|)))}{1-\gamma}$
It is noteworthy that the optimal policy for the optimization formulation (in Lemma 2.2) and the optimal control problem (in Theorem 2.1) are identical. In other words, when $K=K^{*}$ and $\Sigma=\Sigma^{*}$ , the problem attains its optimal solution and $P_{K}$ satisfy the ARE in (2.1). The following lemma provides the explicit form of the first-order derivative of cost function with respect to $K$ and $\Sigma$.
Lemma 2.3 (Explicit form of $\nabla_{K}f(K,\Sigma)$ and $\nabla_{\Sigma}f(K,\Sigma)$).
Assume that $\gamma<1$, it holds that
$\nabla_{K}f(K,\Sigma)=E_{K}S_{K},$
$\nabla_{\Sigma}f(K,\Sigma)=(1-\gamma)^{-1}\left((R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{\top}-\frac{\tau}{2}\Sigma^{-1}\right),$
where $E_{K}=2RK+2\gamma P_{K}[(B^{\top}B+D^{\top}D)K-AB^{\top}]$, $S_{K}=\sum_{t=0}^{\infty}\mathbb{E}x_{t}^{2}$.
3 Global Convergence of Regularized Policy Gradient
It is natural to utilize a gradient descent method for addressing the optimization problem. In [23], the Regularized Policy Gradient (RPG) algorithm was introduced and shown to achieve global optimality in the context of noisy linear quadratic problems. In this chapter, we extend these results by demonstrating that RPG remains globally optimal for stochastic linear quadratic problems with multiplicative noise.
Consider RPG with following updating rules with a fixed step size $\eta_{1}$ and $\eta_{2}$:
$K$ $\leftarrow K-\eta_{1}\frac{\nabla_{K}f(K,\Sigma)}{S_{K,\Sigma}},$ (12)
$\Sigma$ $\leftarrow\Sigma-\eta_{2}\Sigma\nabla_{\Sigma}f(K,\Sigma)\Sigma.$ (13)
By Lemma 2.3, the above update can be written as
$K$ $\leftarrow K-\eta_{1}E_{K},$
$\Sigma$ $\leftarrow\Sigma-\frac{\eta_{2}}{1-\gamma}\Sigma\left((R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{\top}-\frac{\tau}{2}\Sigma^{-1}\right)\Sigma.$
Before proving Global Convergence of Gradient Methods, we first introduce the following lemmas, which establish the gradient dominance condition and the smoothness property of the value function, both of which play a crucial role in the subsequent proofs.
Lemma 3.1 (Gradient Domination of $f(K,\Sigma)$).
Let $(K^{*},\Sigma^{*})\in\Omega$ be an global optimal policy. Assume that $(K,\Sigma)\in\Omega$ and $\mu>0$. Then we have
$\lambda_{1}E_{K}^{\top}E_{K}\leq f(K,\Sigma)-f(K^{*},\Sigma^{*})\leq\lambda_{2}\nabla_{K}f^{\top}(K,\Sigma)\nabla_{K}f(K,\Sigma)+\frac{(1-\gamma)\text{Tr}[(\nabla_{\Sigma}C_{K,\Sigma}(x_{0}))^{2}]}{\sigma_{min}(R)}$ (14)
where $\lambda_{1}=\frac{\mu}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ and $\lambda_{2}=\frac{1}{\mu\sigma_{min}(R)}$
Lemma 3.2 (Gradient Norm Bounds).
The gradient of $f(K,\Sigma)$ have the following bounds,
$\|\nabla_{K}f(K,\Sigma)\|\leq\overline{\|\nabla_{K}f(K,\Sigma)\|}:=\frac{f(K,\Sigma)}{Q}\sqrt{\frac{f(K,\Sigma)-f(K^{*},\Sigma^{*})}{\lambda_{1}}}$
and
$\|\nabla_{\Sigma}f(K,\Sigma)\|\leq\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}:=(1-\gamma)^{-1}\big[\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2\sigma_{min}(\Sigma)}\big]$
Now we have proved $f(K,\Sigma)$ is gradient dominated. If $f(K,\Sigma)$ is smooth and gradient dominated, then the gradient descent methods will convergence to the global optimal at a linear rate. Unfortunately, $f(K,\Sigma)$ cannot satisfy the smoothness condition; this is due to $f(K,\Sigma)=\infty$ when $\gamma V_{K}\geq 1$. We consider the case where the policy $(K,\Sigma)$ is not too close to the boundary, the objective satisfies an almost smoothness condition as follows:
Lemma 3.3 (”Almost” smoothness of $f(K,\Sigma)$).
Fix $0<a<1$ and define $m=\frac{log(a)-a+1}{(a-1)^{2}}$, any $\Sigma$ and $\Sigma^{\prime}$ satisfies $aI\prec\Sigma\prec I$ and $aI\prec\Sigma^{\prime}\prec I$, we have,
$f(K^{\prime},\Sigma^{\prime})-f(K,\Sigma)$ $=S_{K}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
$\quad+q_{K,\Sigma^{\prime}}-q_{K,\Sigma}$
$\leq S_{K}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
$\quad+\frac{Tr\left(((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}\Sigma^{-1})(\Sigma^{\prime}-\Sigma)\right)}{(1-\gamma)}+\frac{\tau m}{2(1-\gamma)}Tr((\Sigma^{-1}\Sigma^{\prime}-I)^{2})$
From the above lemmas, we have established that $f(K,\Sigma)$ possesses the properties of Gradient Domination and is ”almost” smooth. These properties make it possible to prove global convergence. Now we need to show that one step update guarantees a decrease in $f(K,\Sigma)$. To this end, we first prove that the update of $\Sigma$ is bounded.
Lemma 3.4 (Boundedness of update $\Sigma$).
Let $(K,\Sigma)\in\Omega$ be given such that $0\prec\Sigma\preceq I$. Assume $\tau\in(0,2\sigma_{min}(R))$. Fix $a\in(0,\text{min}\{\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|},\sigma_{min}(\Sigma)\})$ with $\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}$. Update of $\Sigma$ will have $aI\prec\Sigma^{\prime}\prec I$.
The boundedness of the update to $\Sigma$ ensures that the cost function remains well-defined along the trajectory during the execution of RPG. Furthermore, we need to show that one step update guarantees a decrease in $f(K,\Sigma)$.
Lemma 3.5 (Contraction of RPG).
Let $(K,\Sigma)\in\Omega$ be given such that $0\prec\Sigma\preceq I$. Assume $\tau\in(0,2\sigma_{min}(R))$. Fix $a\in(0,\text{min}\{\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|},\sigma_{min}(\Sigma)\})$. For $\eta_{1}\leq\frac{1}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ and $\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}$, and $0<\phi=min\{\eta_{1}\mu\frac{\sigma_{min}{(R)}}{S_{K^{*},\Sigma^{*}}},\frac{\eta_{2}\sigma_{min}(R)}{2(1-\gamma)}\}<1$
$f(K^{\prime},\Sigma^{\prime})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{*},\Sigma^{*}))$
Lemma 3.6 (Lower bound of $f(K,\Sigma)$).
For any $(K,\Sigma)\in\Omega$, $f(K,\Sigma)$ has the following lower bound:
$f(K,\Sigma)\geq\mu P_{K}+\frac{\tau k}{2(1-\gamma)}log(\frac{\sigma_{min}(R)}{\pi\tau}).$
With the above lemmas established, we are now ready to prove the following theorem.
Theorem 3.1 (Global convergence of RPG).
Given $\tau\in(0,2\sigma(R)]$, $\epsilon\in(0,1)$ take $(K,\Sigma)\in\Omega$ such that $\Sigma\preceq I$. For
$\eta_{1}=\min\bigg\{\frac{1}{R+\frac{\gamma}{\mu}\|B^{\top}B+D^{\top}D\|\left(f(K)-\frac{\tau K}{2(1-\gamma)}log(\frac{\sigma_{min}(R)}{\pi\tau})\right)},\frac{2}{\tau\sigma_{min}(\Sigma)}\bigg\},$
$\eta_{2}=2\tau(1-\gamma)\eta_{1}^{2}$, and for
$N\geq\max\left\{\frac{\|S_{K^{*},\Sigma^{*}}\|}{2\mu\eta_{1}\sigma_{min}(R)},\frac{1}{\tau^{2}\eta_{1}^{3}\sigma_{min}(R)}\right\}\log\frac{f(K,\Sigma)-f(K^{*},\Sigma^{*})}{\epsilon},$
the Regularized Policy Gradient (RPG) has the following performance bound:
$f(K^{(N)},\Sigma^{(N)})-f(K^{*},\Sigma^{*})\leq\epsilon.$
Proof.
From lemma 3.6 we have
$\frac{1}{R+\gamma P_{K}(B^{\top}B+D^{\top}D)}$ $\geq\frac{1}{\|R\|+\gamma P_{K}\|B^{\top}B+D^{\top}D\|}$
$\geq\frac{1}{R+\frac{\gamma}{\mu}\|B^{\top}B+D^{\top}D\|\left(f(K)-\frac{\tau k}{2(1-\gamma)}log(\frac{\sigma_{min}(R)}{\pi\tau})\right)}$
$\geq\eta_{1}.$
Define $a=\tau\eta_{1}\leq\frac{\tau}{\|R\|+\gamma P_{K}\|B^{\top}B+D^{\top}D\|}$. We will prove this theorem by induction. At $t=0$, we have $\eta_{1}\leq\frac{1}{R+\gamma P_{K}(B^{\top}B+D^{\top}D)},\eta_{2}=2\tau(1-\gamma)\eta_{1}^{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}$. Then we can apply lemma 3.5 such that
$f(K^{(1)},\Sigma^{(1)})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{*},\Sigma^{*})),$
and $aI\preceq\Sigma^{(1)}\preceq I$, where $\phi$ is defined in Lemma 3.5.
Assume the theorem holds at time $t$, then we have $f(K^{(n)},\Sigma^{(n)})\leq f(K^{(n-1)},\Sigma^{(n-1)})\leq f(K,\Sigma)$, and $aI\prec\Sigma^{(t)}\prec I$. Then we have
$\eta_{1}$ $\leq\frac{1}{R+\frac{\gamma}{\mu}\|B^{\top}B+D^{\top}D\|\left(f(K,\Sigma)-\frac{\tau k}{2(1-\gamma)}log(\frac{\sigma_{min}(R)}{\pi\tau})\right)}$
$\leq\frac{1}{R+\frac{\gamma}{\mu}\|B^{\top}B+D^{\top}D\|\left(f(K^{(n)},\Sigma^{(n)})-\frac{\tau k}{2(1-\gamma)}log(\frac{\sigma_{min}(R)}{\pi\tau})\right)},$
and $a\leq\frac{\tau}{\|R\|+\gamma P_{K^{(n)}}\|B^{\top}B+D^{\top}D\|}$. Now Lemma 3.5 can be applied such that
$f(K^{(n+1)},\Sigma^{(n+1)})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{(n)},\Sigma^{(n)})).$
and $aI\preceq\Sigma^{(n+1)}\preceq I$ The induction is complete. Finally, observe that $0<\phi\leq\frac{2\mu\eta_{1}\sigma_{min}(R)}{\|S_{K^{*},\Sigma^{*}}\|}<1$ and $\phi\leq\frac{\eta_{2}a\sigma_{min}(R)}{2(1-\gamma)}=\tau^{2}\eta_{1}^{3}\sigma_{min}(R)$. The proof is completed.
4 Global Convergence of Sample Based Regularized Policy Gradient
In this section, we consider a model in which all parameters, $A,B,C,D,Q,R$, as well as the exact value of $f(K,\Sigma)$, are unknown. The only available information pertains to the form of the system dynamics and approximate values of the cost trajectories (i.e., $\sum_{t=0}^{l-1}\gamma^{t}(Qx_{t}^{2}+u_{t}^{T}Ru_{t}+\tau\log\pi(u_{t}|x_{t}))$, where $l<\infty$ serves as the rollout length in the simulation environment). Employing a zero-order optimization techniques, we propose the Sample Based Regularized Policy Gradient (SB-RPG) method for our stochastic optimal control problems. This section demonstrates that, even under settings with unknown parameters, our approach achieves globally optimal solutions with high probability. The pseudocode for SB-RPG is provided in Algorithm 1, where $\widehat{\nabla}_{K},\widehat{\nabla}_{\Sigma},$ and $\widehat{S}$ denote sample based estimate of $\nabla_{K}f(K,\Sigma),\nabla_{\Sigma}f(K,\Sigma)$ and $S_{K,\Sigma}$, respectively.
Algorithm 1 Pseudocode code of Sample Based Regularized Policy Gradient (SB-RPG)
Input: initial policy $(K,\Sigma)\in\Omega$, updating steps NN, policy estimate trajectories $M$, roll out length $l$, smoothing parameters $r_{1}$ and $r_{2}$.
for $j=1,\cdots N$ do
  for $i=1,\cdots M$ do
   Sample a policy $K_{i}=K+U_{i}$, where $U_{i}$ is drawn uniformly at random over $\|U_{i}\|_{F}=r_{1}$.
   Simulate $f^{(l)}_{i}(K_{i},\Sigma)=\sum_{t=0}^{\ell-1}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{T}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)$ and $S_{i}^{(l)}=\sum_{t=0}^{l-1}\gamma^{t}x_{t}^{2}$ under policy $(K_{i},\Sigma)$ for $l$ steps.
  end for
  Estimate: $\widehat{\nabla}_{K}=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r_{1}^{2}}f^{(l)}_{i}(K,\Sigma)U_{i}$, $\widehat{S}=\frac{1}{M}\sum_{i=1}^{M}S_{i}^{(l)}$
  Update: $K\leftarrow K-\eta_{1}{\widehat{\nabla}_{K}}/{\widehat{S}}$
  for $i=1,\cdots M$ do
   Sample a policy $\Sigma_{i}=\Sigma+V_{i}$, where $V_{i}$ is drawn uniformly at random over $\|V_{i}\|_{F}=r_{2}$.
   Simulate the cost of $f^{(l)}_{i}(K,\Sigma_{i})=\sum_{t=0}^{l-1}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{T}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)$ under policy $(K,\Sigma_{i})$ for $l$ steps
  end for
  Estimate: $\widehat{\nabla}_{\Sigma}=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r_{2}^{2}}f^{(l)}_{i}(K,\Sigma_{i})V_{i}$
  Update: $\Sigma\leftarrow\Sigma-\eta_{2}\Sigma\widehat{\nabla}_{\Sigma}\Sigma$
end for
To prove global convergence of SB-RPG, we need to prove step by step that all sample-based estimates, under some condition on $l,M$ and $r$, can be $\epsilon$ close to the true value with high probability. To this end, perturbation analysis and several other technical tools are essential. We present the following lemmas, with proofs provided in the appendix.
Lemma 4.1.
For any $(K,\Sigma)\in\Omega$, $S_{K,\Sigma}$ can be written as
$S_{K,\Sigma}=\mu\sum_{t=0}^{\infty}(\gamma V_{k})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{1}{1-\gamma}-\frac{1}{1-\gamma V_{K}}\right].$
Furthermore, $S_{K,\Sigma}$ has the following bound
$\frac{\mu}{1-\gamma V_{K}}\leq S_{K}\leq\frac{f(K,\Sigma)-{(1-\gamma)^{-1}}\left[Tr(\Sigma R)-\frac{\tau}{2}\left(n+log(2\pi)^{n}|\Sigma|\right)\right]}{Q}.$
Lemma 4.2 (Approximate $f(K,\Sigma)$ and $S_{K,\Sigma}$ with any desired accuracy).
For any $K,\Sigma$ with $f(K,\Sigma)<\infty$, let $f^{(l)}(K,\Sigma)=\mathbb{E}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{T}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right)\right]$ and $S^{(l)}_{K,\Sigma}=\sum_{t=0}^{l-1}\gamma^{t}\mathbb{E}x_{t}^{2}$. we have
(i) $S_{K,\Sigma}-S^{(l)}_{K,\Sigma}\leq\epsilon$, if
$l\geq\frac{\log\epsilon-\log S_{K,\Sigma}}{log\gamma}$
(ii) $f(K,\Sigma)-f^{(l)}(K,\Sigma)\leq\epsilon$, if
$l\geq\frac{\log\epsilon-\log\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]}{log\gamma}$
Lemma 4.3 ($S_{K,\Sigma}$ Perturbation).
If $\|K-K^{\prime}\|\leq h_{\Sigma}$ and $\|\Sigma^{\prime}-\Sigma\|\leq\|\Sigma\|$, then
$|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|\leq h_{K}\|K-K^{\prime}\|+h_{2}\|\Sigma-\Sigma^{\prime}\|,$
where
$h_{\Sigma}$ $=\frac{1}{2S_{K,\Sigma}^{2}}\frac{\mu^{2}}{\sqrt{\frac{1}{2}\frac{\mu^{2}\|B^{\top}B+D^{\top}D\|}{S_{K,\Sigma}^{2}}+\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|^{2}}+\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|},$
$h_{2}$ $=\frac{\gamma Tr((B^{\top}B+D^{\top}D))}{(1-\gamma)(1-\gamma V_{K})},$
$g_{\Sigma}$ $=2\left(\frac{1}{1-\gamma V_{K}}\right)^{2}(2\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|+\|B^{\top}B+D^{\top}D\|h_{\Sigma}),$
$h_{K}$ $=2g_{\Sigma}\frac{(1-\gamma)\mu+\gamma Tr(\Sigma(B^{\top}B+D^{\top}D))}{(1-\gamma)}.$
Lemma 4.4 ($P_{K}$ perturbation).
If $\|K^{\prime}-K\|\leq\min\{h_{\Sigma},\|K\|\}$, then
$|P_{K^{\prime}}-P_{K}|\leq h_{5}\|K^{\prime}-K\|$
where
$h_{5}=\frac{3\|K\|\|R\|}{1-\gamma V_{K}}+(Q+4\|R\|\|K\|^{2})g_{\Sigma}.$
Lemma 4.5 ($\nabla_{K}f(K,\Sigma)$ and $\nabla_{\Sigma}f(K,\Sigma)$ perturbation).
If $\|K^{\prime}-K\|\leq\min\{h_{\Sigma},\|K\|\}$ and $\|\Sigma^{\prime}-\Sigma\|_{F}\leq\min\{\frac{\sigma_{min}(\Sigma)}{2},\|\Sigma\|\}$, then
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|\leq h_{6}\|K^{\prime}-K\|+h_{7}\|\Sigma^{\prime}-\Sigma\|$
and
$\|\nabla_{\Sigma}f(K^{\prime},\Sigma^{\prime})-\nabla_{\Sigma}f(K,\Sigma)\|\leq h_{8}\|K^{\prime}-K\|+h_{9}\|\Sigma^{\prime}-\Sigma\|$
where
$h_{E}$ $=2(\|R\|+\gamma A\cdot h_{5}\|B\|+\gamma P_{K}(B^{\top}B+D^{\top}D)+2\gamma\cdot h_{5}\|B^{\top}B+D^{\top}D\|\|K\|),$
$h_{6}$ $=h_{K}\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}+h_{E}|S_{K^{\prime},\Sigma^{\prime}}|,\quad h_{9}=\frac{\tau\sigma_{min}(\Sigma)}{4(1-\gamma)},$
$h_{7}$ $=h_{2}\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|},\quad h_{8}=\frac{\gamma(B^{\top}B+D^{\top}D)}{(1-\gamma)}h_{5}.$
We define $f_{r_{1}}(K,\Sigma):=\mathbb{E}_{U\sim\mathbb{B}_{r_{1}}}[f(K+U,\Sigma)],$ and $f_{r_{2}}(K,\Sigma):=\mathbb{E}_{V\sim\mathbb{B}_{r_{2}}}[f(K,\Sigma+V)]$. where $\mathbb{B}_{r}$ denotes the uniform distribution over the points with norm $r$ (boundary of a sphere). The following lemma shows that the gradient of $f_{r_{1}}(K,\Sigma)$ and $f_{r_{2}}(K,\Sigma)$ can be estimated with an oracle for the function value.
Lemma 4.6.
$\nabla_{K}f_{r_{1}}(K,\Sigma)=\frac{n}{r_{1}^{2}}\mathbb{E}_{U\sim\mathbb{S}_{r_{1}}}[f(K+U,\Sigma)U].$
and
$\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)=\frac{n^{2}}{r_{2}^{2}}\mathbb{E}_{V\sim\mathbb{S}_{r_{2}}}[f(K,\Sigma+V)V].$
where $\mathbb{S}_{r}$ denotes the uniform distribution over all points with norm at most r (the entire sphere).
The following lemmas show that $\nabla_{K}f(K,\Sigma)$, $\nabla_{\Sigma}f(K,\Sigma)$ and $S_{K,\Sigma}$ can be estimated with finite samples under small perturbation at any desired accuracy.
Lemma 4.7 (Estimate of $\nabla_{K}f(K,\Sigma)$).
Given an arbitrary tolerance $\epsilon_{1}>0$ and probability $\kappa_{1}\in(0,1)$. Let $x_{t}^{i},u_{t}^{i}$ be $i$-th single path sampled using policy $(K+U_{i},\Sigma)\in\Omega$, where $\|U_{i}\|_{F}\leq r_{1}$, define
$\widehat{\nabla}_{K}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r_{1}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]U_{i}.$
Assume that (i) the distribution of the initial states implies that $\|x_{0}^{i}\leq L\|$ almost surely for any $i$. (ii) the multiplicative noises are distributed such that $\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$ for any $i$. If set
$r_{1}$ $\leq\frac{\epsilon_{1}}{2h_{6}}$
$M$ $\geq\max\big\{\frac{2n}{(\epsilon_{1}/6)^{2}}(\sigma_{1}+\frac{R_{1}\epsilon}{18\sqrt{n}})\log\big(\frac{n+1}{\sqrt{\kappa_{1}}}\big),\frac{2n}{(\epsilon/3)^{2}}(\sigma^{2}_{2}+\frac{R_{2}\epsilon}{9x\sqrt{n}})\log(\frac{n+1}{\sqrt{\kappa_{1}}})\big\}$
$l$ $\geq\log(\gamma)^{-1}\left[\log\left(\frac{r_{1}}{n}\cdot\frac{\epsilon}{3}\right)-\log\left(2|f(K,\Sigma)|\left(2\|K\|^{2}\|R\|+\frac{1}{|Q|}\right)+|\psi|\left(1+\frac{1}{|Q|}+\frac{1}{1-\gamma}\right)\right)\right]$
then
$\widehat{\nabla}_{K}-f(K,\Sigma)||_{F}\leq\epsilon$
with high probability (at least $1-\kappa_{1}$), where $\psi_{1}=Tr(\Sigma R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma|),\sigma_{1}=\big(\frac{2n}{r_{1}}f(K,\Sigma)\big)^{2}+\big(\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}\big)^{2},R_{1}=\frac{2n}{r_{1}}f(K,\Sigma)+\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|},\sigma_{2}=(2\Gamma L^{2}f(K,\Sigma)r_{1})^{2}+(\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|})^{2},R_{2}=2\Gamma L^{2}f(K,\Sigma)r_{1}+\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$.
Lemma 4.8 (Estimate of $\nabla_{\Sigma}f(K,\Sigma)$).
Given an arbitrary tolerance $\epsilon_{2}>0$ and probability $\kappa_{2}\in(0,1)$. Let $x_{t}^{i},u_{t}^{i}$ be $i$-th single path sampled using policy $(K,\Sigma+V_{i})\in\Omega$, where $\|V_{i}\|_{F}\leq r_{2}$, define
$\widehat{\nabla}_{\Sigma}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r_{2}^{2}}\left[\sum_{t=0}^{l-1}\gamma^{t}\left(Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}Ru_{t}^{i}+\tau\log\pi(u_{t}|x_{t})\right)\right]V_{i}.$
If set
$r_{2}$ $\leq\frac{\epsilon_{2}}{2h_{9}}$
$M$ $\geq\max\big\{\log\left(\frac{n}{\kappa_{2}}\right)\cdot\frac{4}{\epsilon_{2}^{2}}\left(2r_{2}^{2}+R_{2}\frac{\epsilon_{2}}{3}\right),\frac{6}{\epsilon_{2}^{2}}\log\left(\frac{n}{\kappa_{2}}\right)\left(3r_{2}^{2}+R_{2}^{\prime}\frac{\epsilon_{2}}{3}\right)\big\}$
$l$ $\geq\frac{\log\left(\frac{\epsilon_{2}r_{2}}{3n}\right){\log(\gamma)}-\log\left[\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}\right)2f(K,\Sigma)+\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\right)|\psi_{2}|\right]}{\log(\gamma)}$
then
$\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)||_{F}\leq\epsilon_{2}$
with high probability (at least $1-\kappa_{2}$), where $\psi_{2}=\text{Tr}(\Sigma R)+\frac{\tau}{2}(n+2\log(2\pi)^{n}|\Sigma|)$ , $R_{2}=\frac{2nf(K,\Sigma)}{r_{2}}+\frac{\epsilon_{2}}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$.
Lemma 4.9 (Estimate of $S_{K,\Sigma}$ under perturbation).
Given an arbitrary tolerance $\epsilon_{3}>0$ and probability $\kappa_{3}\in(0,1)$. Let $x_{t}^{i},u_{t}^{i}$ be a single path sampled using policy $(K+U_{i},\Sigma)\in\Omega$, where $\|U_{i}\|_{F}\leq r_{3}$. Define
$\widehat{S}_{K,\Sigma}:=\frac{1}{M}\sum_{i=1}^{M}\sum_{t=0}^{l-1}\gamma^{t}(x_{t}^{i})^{2}.$
If set $r_{3}\leq\min\{\frac{S_{K,\Sigma}}{2h_{K}},\frac{\epsilon_{3}}{3h_{K}},h_{\Sigma}\}$, $l\geq\frac{\log\epsilon_{3}/3-\log S_{K,\Sigma}/2}{log\gamma}$ and $M\geq\sqrt{\frac{3S_{K,\Sigma}}{\epsilon_{3}}\log\frac{n}{\kappa_{3}}}$, then
$|S_{K,\Sigma}-\widehat{S}_{K,\Sigma}|<\epsilon$
with high probability (at least $1-\kappa_{3}$). Furthermore, if $\epsilon_{3}\leq\mu/2$, then $\widehat{S}_{K,\Sigma}\geq\mu/2$ .
Lemma 4.10 ($f({K,\Sigma})$ perturbation).
We have
$|f(K,\Sigma)-f(K^{\prime},\Sigma^{\prime})|\leq h_{10}\|K^{\prime}-K\|+h_{11}\|\Sigma^{\prime}-\Sigma\|$
if $\|K^{\prime}-K\|\leq\min\{h_{\Sigma},\|K\|\}$ and $\|\Sigma^{\prime}-\Sigma\|\leq\|\Sigma\|$, where $h_{10}=({2\gamma\|\Sigma\|\|B^{\top}B+D^{\top}D\|}(1-\gamma)^{-1}+\mu)h_{5}$ and $h_{11}=\frac{m\|\Sigma^{-1}\|_{F}}{2}+\overline{\|\nabla_{\Sigma}q_{K,\Sigma}\|}.$
With the above lemmas, we can now prove the following theorem.
Theorem 4.1 (Global Convergence of SB-RPG).
Given an arbitrary tolerance $\epsilon>0$ and probability $\kappa\in(0,1)$. If we set
(i) $\eta_{1},\eta_{2}$, $\phi$ to be equal to the values in Theorem 3.1.
(ii) $N\geq N_{SB}$, where $N_{SB}=\frac{N_{RPG}\log(1-\phi)}{\log(1-{\phi}/{2})}$ and $N_{RPG}$ denotes the minium update steps in Theorem 3.1.
(iii) $M$, $l$, $r_{1}$ and $r_{2}$ satisfy the conditions in Lemma 4.7, 4.8, and 4.9 when $\kappa_{1}=\kappa_{3}=1-(1-\kappa)^{1/(4N_{SB})}$, $\kappa_{2}=1-(1-\kappa)^{1/(2N_{SB})}$, $\epsilon_{1}=\frac{\mu\phi\epsilon}{8\eta_{1}(h_{10}+h_{11})}$, $\epsilon_{2}=\frac{\phi\|\Sigma\|\epsilon}{2\eta_{2}(h_{10}+h_{11})}$ and $\epsilon_{3}=\frac{\mu^{2}\phi\epsilon}{8\eta_{1}\overline{\|{\nabla_{K}f(K,\Sigma)}\|}(h_{10}+h_{11})}$. Then SB-RPG (in Algorithm 1) will have the following performance bound after NN times update
$f(K^{(N)},\Sigma^{(N)})-f(K^{*},\Sigma^{*})<\epsilon$
with high probability (at least $1-\kappa$).
Proof.
Define $K^{\prime},\Sigma^{\prime}$ as the result of one step update of RPG in (12) and (13). In Lemma 3.5 we have when $\eta_{1}$ and $\eta_{2}$ are chosen properly, we have
$f(K^{\prime},\Sigma^{\prime})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{*},\Sigma^{*})),$
Define $K^{\prime\prime}=K-\eta_{1}\frac{\widehat{\nabla}_{K}}{\widehat{S}_{K,\Sigma}}$ and $\Sigma^{\prime\prime}=\Sigma-\eta_{2}\Sigma\widehat{\nabla}_{\Sigma}\Sigma$ where $\widehat{\nabla}_{K}$ and $\widehat{\nabla}_{\Sigma}$ are defined in Lemma 4.7 and Lemma 4.8. We will show that when $\nabla_{K}f(K,\Sigma),\nabla_{\Sigma}f(K,\Sigma)$ and $S_{K,\Sigma}$ are estimated accurately enough, then we have $|f(K^{\prime\prime},\Sigma^{\prime\prime})-f(K^{\prime},\Sigma^{\prime})|\leq\frac{\epsilon}{2}\phi$ , which implies that when $f(K^{\prime},\Sigma^{\prime})-f(K^{*},\Sigma^{*})>\epsilon$,
$f(K^{\prime\prime},\Sigma^{\prime\prime})-f(K^{*},\Sigma^{*})\leq(1-\frac{\phi}{2})(f(K,\Sigma)-f(K^{*},\Sigma^{*}))$
with probability $(1-\kappa)^{1/N_{SB}}$. As we proved perturbation of $f(K,\Sigma)$ in Lemma 4.10, we only need to establish the following two claims, both under condition given in the theorem.
(i) $\|K^{\prime\prime}-K^{\prime}\|\leq\frac{\phi\epsilon}{2(h_{10}+h_{11})}$ with probability $(1-\kappa)^{1/(2N_{SB})}$.
$\|K^{\prime\prime}-K^{\prime}\|$ $=\eta_{1}\Bigg\|\frac{\widehat{\nabla}_{K}}{\widehat{S}_{K,\Sigma}}-\frac{\nabla_{K}f(K,\Sigma)}{S_{K,\Sigma}}\Bigg\|$
$\leq\eta_{1}\frac{1}{\widehat{S}_{K,\Sigma}}\|\nabla_{K}f(K,\Sigma)-\widehat{\nabla}_{K}\|+\eta_{1}\overline{\|{\nabla_{K}f(K,\Sigma)}\|}\Bigg|\frac{1}{S_{K,\Sigma}}-\frac{1}{\widehat{S}_{K,\Sigma}}\Bigg|$
For the first term, from Lemma 4.9 and Lemma 4.7 we have
$\frac{\eta_{1}}{\widehat{S}_{K,\Sigma}}\|\nabla_{K}f(K,\Sigma)-\widehat{\nabla}_{K}\|\leq\frac{2\eta_{1}}{\mu}\|\nabla_{K}f(K,\Sigma)-\widehat{\nabla}_{K}\|\leq\frac{\epsilon\phi}{4(h_{10}+h_{11})}$
with probability at least $(1-\kappa)^{1/(4N_{SB})}$, as we set $\epsilon_{1}=\frac{\mu\phi}{8\eta_{1}(h_{10}+h_{11})}\epsilon$ and $1-\kappa_{1}=(1-\kappa)^{1/(4N_{SB})}$. For the second term, by standard matrix perturbation and Lemma 4.9, we have
$\eta_{1}\|{\nabla_{K}f(K,\Sigma)}\|\Bigg|\frac{1}{S_{K,\Sigma}}-\frac{1}{\widehat{S}_{K,\Sigma}}\Bigg|\leq\eta_{1}\overline{\|{\nabla_{K}f(K,\Sigma)}\|}\frac{2|\widehat{S}_{K,\Sigma}-S_{K,\Sigma}|}{\mu^{2}}\leq\frac{\epsilon\phi}{4(h_{10}+h_{11})}$
with probability $(1-\kappa)^{1/(4N)}$, as we set $\epsilon_{3}=\frac{\mu^{2}\phi}{8\eta_{1}\overline{\|{\nabla_{K}f(K,\Sigma)}\|}(h_{10}+h_{11})}\epsilon$ and $1-\kappa_{3}=(1-\kappa)^{1/(4N_{SB})}$.
(ii) $\|\Sigma^{\prime\prime}-\Sigma^{\prime}\|\leq\frac{\phi\epsilon}{2(h_{10}+h_{11})}$ with probability $(1-\kappa)^{1/2N_{SB}}$.
$\|\Sigma^{\prime\prime}-\Sigma^{\prime}\|$ $=\eta_{2}\|\Sigma\nabla_{\Sigma}f(K,\Sigma)\Sigma-\Sigma\widehat{\nabla}_{\Sigma}\Sigma\|$
$\leq\eta_{2}\|\Sigma\|\|\nabla_{\Sigma}f(K,\Sigma)\Sigma-\widehat{\nabla}_{\Sigma}\Sigma\|$
$\leq\eta_{2}\|\Sigma\|^{2}\|\nabla_{\Sigma}f(K,\Sigma)-\widehat{\nabla}_{\Sigma}\|$
From Lemma 4.8, we have $\|\Sigma^{\prime\prime}-\Sigma^{\prime}\|\leq\frac{\epsilon\phi}{2(h_{10}+h_{11})}$ with probability $(1-\kappa)^{1/2N}$, as we set $\epsilon_{2}=\frac{\phi}{2\|\Sigma\|^{2}\eta_{2}(h_{10}+h_{11})}\epsilon$ and $1-\kappa_{2}=(1-\kappa)^{1/(2N_{SB})}$ .
Combining (i), (ii) and Lemma 4.10, we have
$|f(K^{\prime\prime},\Sigma^{\prime\prime})-f(K^{\prime},\Sigma^{\prime})|$ $\leq h_{10}\|K^{\prime\prime}-K^{\prime}\|+h_{11}\|\Sigma^{\prime\prime}-\Sigma^{\prime}\|\leq\frac{\epsilon}{2}\phi$
with probability $(1-\kappa)^{1/N_{SB}}$. Then we have
$f(K^{\prime\prime},\Sigma^{\prime\prime})-f(K^{*},\Sigma^{*})\leq\left(1-\frac{\phi}{2}\right)(f(K,\Sigma)-f(K^{*},\Sigma^{*}))$
with probability $(1-\kappa)^{1/N_{SB}}$ for each iteration. Now we have proved contraction of SB-RPG, the rest of the proof remains the same as Theorem 3.1, the probability of convergence becomes $((1-\kappa)^{1/N_{SB}})^{N_{SB}}=1-\kappa$ after NN times iteration. ∎
5 Numerical Experiments
In this section we provide a 3 dimensional control and 1 dimensional state numerical experiment using SB-RPG (Algorithm 1), where all the parameters are unknown. Although SB-RPG does not require the exact values of the parameters, we need to specify the following parameters to simulate the system cost of $f(K,\Sigma)$. We set $A=0.7,B=(0.1,0.2,0.3),C=0.03,Q=0.5,\gamma=0.5,\tau=0.1,$
$D=\begin{pmatrix}0.05&0.13&0.12\\ 0.13&0.07&0.10\\ 0.12&0.10&0.03\end{pmatrix}\quad R=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{pmatrix}$.
The theoretical analysis provides rather conservative limits for the step size $\eta$, number of rollouts $M$, and rollout length $l$. To ensure practicality, we determined the constant step size, number of rollouts, rollout length, and exploration radius by performing a grid search over a set of reasonable values. In simulations, we obtained the baseline optimal cost $f(K^{*},\Sigma^{*})$ by solving the ARE in (2) to high accuracy ($e^{-5}$ accuracy) using value iteration.
(a) Expected cost of RPG
(b) Relative error of $f(K,\Sigma)$ and $f(K^{*},\Sigma^{*})$
Figure 1: Expected cost of RPG
In Figure 1, we compare the optimal cost $f(K^{*},\Sigma^{*})$ with the SB-RPG cost $f(K,\Sigma)$ throughout the training process. The left subfigure shows the absolute error $|f(K,\Sigma)-f(K^{*},\Sigma^{*})|$, while the right subfigure displays the relative error $\frac{|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}{f(K^{*},\Sigma^{*})}$. As observed in the right subfigure, the relative error remains approximately 5%.
(a) Squared error of $K$ and $K^{*}$
(b) Squared error of $\Sigma$ and $\Sigma^{*}$
Figure 2: Squared error between optimal policy and SB-RPG
In Figure 2, we present the relative error between the sample-based policy and the optimal policy. The left subfigure illustrates the squared error $\|K^{\prime}-K\|_{F}^{2}$, and the right subfigure shows the squared error $\|\Sigma^{\prime}-\Sigma\|_{F}^{2}$.
Acknowledgments
I first want to thank Dr. Xun Li for granting me the opportunity to join his research group and contribute to the meaningful work being done with his students. I would also like to extend my thanks to the Applied Math Department at The Hong Kong Polytechnic University for welcoming me during my two-weeks stay in Hong Kong.
Appendix A Proofs in Section 2
A.1 Proof of Lemma 2.1
The proof of Lemma 2.1 is provided in Section 8.1 of [23].
A.2 Proof of Lemma 2.2
Proof.
We use the similar augment with Theorem 2.1. Take the policy $(K,\Sigma)$ into objective function we have
$J(x)$ $=Qx^{2}+\mathbb{E}_{\pi}\Bigl\{u^{T}Ru+\tau\log(\pi(u|x))+\gamma\left[P((A+w^{x}C)x+(B+w^{u}D)u)^{2}+q\right]\Bigr\}$
$=(Q+\gamma P(A^{2}+C^{2}))x^{2}+\gamma q$
$\quad+\mathbb{E}_{\pi}\Bigl\{u^{\top}(R+\gamma P(B^{\top}B+D^{\top}D))u+\tau\log(\pi(u|x))+2\gamma APxBu\Bigr\}.$
$=(Q+K^{\top}RK+\gamma P_{K}(A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK))x^{2}$
$\quad+\gamma q_{K,\Sigma}-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|))+\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))).$
Take the above equals to $P_{K}x^{2}+q_{K,\Sigma}$, we have:
$P_{K}=Q+K^{\top}RK+\gamma P_{K}(A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK),$
$q_{K,\Sigma}=\frac{\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|)))}{1-\gamma}$
A.3 Proof of Lemma 2.3
Proof.
Define $V_{K}=A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK$. We know $f(K,\Sigma)=\mathbb{E}_{x\sim\mathcal{D}}[C(K,\Sigma)]=P_{K}\mathbb{E}x^{2}+q_{K,\Sigma}$. $\nabla_{\Sigma}f(K,\Sigma)$ can be computed directly:
$\nabla_{\Sigma}f(K,\Sigma)$ $=\nabla_{\Sigma}\left(q_{K,\Sigma}\right)$
$=\nabla_{\Sigma}\left(\frac{\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|)))}{1-\gamma}\right)$
$=(1-\gamma)^{-1}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{\top}-\frac{\tau}{2}\Sigma^{-1}.$
From (1) we have
$\mathbb{E}(x_{t+1}^{2})=V_{K}\mathbb{E}x^{2}_{t}+\text{Tr}(\Sigma(B^{\top}B+D^{\top}D)).$
From (11) we have
$\nabla_{K}P_{K}$ $=2RK+\gamma P_{K}\nabla V_{K}+\gamma\nabla P_{K}V_{K}$
$=(2RK+\gamma P_{K}\nabla V_{K})\sum_{t=0}^{\infty}(\gamma V_{K})^{t}$
and
$\nabla_{K}q_{K,\Sigma}=\sum_{t=0}^{\infty}\gamma^{t+1}\text{Tr}(\Sigma(B^{\top}B+D^{\top}D))\nabla_{K}P_{K}.$
Combining the above we have
$\nabla_{K}f({K,\Sigma})$ $=(2RK+\gamma P_{K}\nabla V_{K})\mathbb{E}x_{0}^{2}+\gamma\nabla_{K}P_{K}V_{K}\mathbb{E}x_{0}^{2})+\sum_{t=0}^{\infty}\gamma^{t+1}\text{Tr}(\Sigma(B^{\top}B+D^{\top}D))\nabla_{K}P_{K}$
$=(2RK+\gamma P_{K}\nabla V_{K})\mathbb{E}x_{0}^{2}+\gamma(V_{K}\mathbb{E}x_{0}^{2}+\text{Tr}(\Sigma(B^{\top}B+D^{\top}D))\nabla P_{K}$
$\quad+\sum_{t=1}^{\infty}\gamma^{t+1}\text{Tr}(\Sigma(B^{\top}B+D^{\top}D))\nabla P_{K}$
$=(2RK+\gamma P_{K}\nabla V_{K})\mathbb{E}x_{0}^{2}+\gamma\mathbb{E}x_{1}^{2}\nabla_{K}P_{K}+\sum_{t=1}^{\infty}\gamma^{t+1}\text{Tr}(\Sigma(B^{\top}B+D^{\top}D))\nabla_{K}P_{K}$
$=(2RK+\gamma P_{K}\nabla V_{K})(\mathbb{E}x_{0}^{2}+\gamma\mathbb{E}x_{1}^{2})$
$\quad+\gamma(\mathbb{E}x_{1}^{2}\nabla P_{K}+\sum_{t=0}^{\infty}\gamma^{t+1}\text{Tr}(\Sigma(B^{\top}B+D^{\top}D))\nabla P_{K})$
$=E_{K}\sum_{t=0}^{\infty}(\gamma^{t}\mathbb{E}x_{t}^{2})$
$=E_{K}S_{K,\Sigma}$
The proof is completed. ∎
Appendix B Proofs in Section 3
B.1 Proof of Lemma 3.1
Proof.
The proof is divided into the following steps
Definition of advantage For any policy $K,K^{\prime}$ that have the finite cost, denote their trajectories as $x_{t}$ and $x^{\prime}_{t}$ respectively. When $x_{0}=x^{\prime}_{0}$,we have the following,
$C_{K^{\prime},\Sigma^{\prime}}(x_{0})-C_{K,\Sigma}(x_{0})$ $=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})\right]\Biggl\}-C_{K,\Sigma}(x_{0})$ (15)
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})-C_{K,\Sigma}(x^{\prime}_{t})+C_{K,\Sigma}(x^{\prime}_{t})\right]\Biggl\}-C_{K}(x_{0})$ (16)
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})-C_{K,\Sigma}(x^{\prime}_{t})\right]+\sum_{t=1}^{\infty}\gamma^{t}C_{K}(x^{\prime}_{t})\Biggl\}$ (17)
$=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})+\gamma C_{K,\Sigma}(x^{\prime}_{t+1})-C_{K,\Sigma}(x^{\prime}_{t})\right]\Biggl\}$ (18)
$\triangleq\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x^{\prime}_{t},u_{t}^{\prime})\Biggl\}$ (19)
where $A_{K}(x,K^{\prime})$ is called ”advantage”, which can be viewed as the change in cost starting at state $x$ between one if choose $u_{t}=-K^{\prime}x$ only at current time and then $u_{t}=-Kx_{t}$ for all $t$ after the current time (i.e., $x^{2}(Q+K^{\prime\top}RK^{\prime})+C_{K}(x_{t+1})$) and one if $u_{t}=-Kx_{t}$ for all $t$ (i.e., $C_{K}(x)$)
We now want to find the $\mathbb{E}_{\pi^{\prime}}\left[A_{K,\Sigma}(x,u^{\prime})\right]$
Expectation
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $=\mathbb{E}_{\pi^{\prime}}\Biggl\{\sum_{t=0}^{\infty}\gamma^{t}\left[(x^{\prime}_{t})^{2}(Q+K^{\prime\top}RK^{\prime})+\tau\log\pi^{\prime}(u_{t}^{\prime}|x^{\prime}_{t})\right]\Biggl\}$ (20)
$+\gamma\mathbb{E}\biggl\{P_{K}\left[(A+w^{x}C)x+(B+w^{u}D)u\right]^{2}\biggl\}-P_{K}x^{2}-q_{K,\Sigma}$ (21)
$=(Q+K^{\prime\top}RK^{\prime})x^{2}+Tr(\Sigma^{\prime}R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma^{\prime}|)-P_{k}x^{2}-(1-\gamma)q_{K,\Sigma}$ (22)
$+\gamma\left[P_{K}(A^{2}+C^{2}+K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-ABK^{\prime})x^{2}+Tr(\Sigma^{\prime}P_{K}(B^{\top}B+D^{\top}D))\right]$ (23)
$=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma^{\prime}|)-P_{k}x^{2}-(1-\gamma)q_{K,\Sigma}$ (24)
$+Tr(\Sigma^{\prime}(R+\gamma P_{K}(B^{\top}B+D^{\top}D)))$ (25)
$=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-P_{K}x^{2}+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (26)
Cost Difference
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $=(Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K}^{\prime})x^{2}-P_{K}x^{2}+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (28)
$=x^{2}[Q+K^{\prime\top}RK^{\prime}+\gamma P_{K}V_{K^{\prime}}-(Q+K^{\top}RK+\gamma P_{K}V_{K})]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (29)
$=x^{2}[K^{\prime\top}RK^{\prime}-K^{\top}RK+\gamma P_{K}(V_{K^{\prime}}-V_{K})]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (30)
$=x^{2}[K^{\prime\top}RK^{\prime}-K^{\top}RK-2\gamma P_{K}AB(K^{\prime}-K)$ (31)
$\quad+\gamma P_{K}(K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-K^{\top}(B^{\top}B+D^{\top}D)K)]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (32)
$=x^{2}[(K+K^{\prime}-K)^{\top}R(K+K^{\prime}-K)-K^{\top}RK-2\gamma P_{K}AB(K^{\prime}-K)$
$\quad+\gamma P_{K}((K+K^{\prime}-K)^{\top}(B^{\top}B+D^{\top}D)(K+K^{\prime}-K)$ (33)
$\quad-\gamma P_{K}K^{\top}(B^{\top}B+D^{\top}D)K)]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (34)
$=x^{2}[K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)$ (35)
$\quad+2(K^{\prime}-K)(R+\gamma P_{K}(B^{\top}B+D^{\top}D))K-\gamma P_{K}AB^{\top}]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (36)
$=x^{2}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$ (37)
$+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (38)
as $\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ is in a quadratic form of $K^{\prime}-K$, we have,
$\mathbb{E}_{\pi^{\prime}}\left[A_{K}(x,K^{\prime})\right]$ $\geq-x^{2}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{\prime}})$ (39)
with equality when $K^{\prime}-K=-(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}$
Upper Bound
Let $S_{K}=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[x_{t}^{2}]$ and remember that $\nabla_{K}f(K,\Sigma)=S_{K}E_{K}$
$C_{K,\Sigma}(x_{0})-C_{K^{*},\Sigma^{*}}(x_{0})$ $=-\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K}(x^{*}_{t},\pi^{*})\right]$
$=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}(x^{*}_{t})^{2}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]-(\gamma-1)(q_{K,\Sigma}-q_{K,\Sigma^{*}})$
$=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}(x^{*}_{t})^{2}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+(q_{K,\Sigma}-q_{K,\Sigma^{*}})$
$=S_{K^{*},\Sigma^{*}}E^{\top}_{K}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}+(q_{K,\Sigma^{*}}-q_{K,\Sigma})$
$\leq\frac{S_{K^{*},\Sigma^{*}}}{\sigma_{min}(R)}{E_{K}^{\top}E_{K}}+Tr(\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma-\Sigma^{*}))$
$\leq\frac{S_{K^{*},\Sigma^{*}}}{\mu^{2}\sigma_{min}(R)}\nabla^{\top}_{K}f_{K,\Sigma}(x_{0})\nabla_{K}f_{K,\Sigma}(x_{0})$
$\quad+Tr[\nabla_{\Sigma}C_{K,\Sigma}(x_{0})((R+\gamma P_{K}(B^{\top}B+D^{\top}D)))^{-1}((R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{\top}-\frac{\tau}{2}\Sigma^{-1})\Sigma]$
$\leq\frac{S_{K^{*},\Sigma^{*}}}{\mu^{2}\sigma_{min}(R)}\nabla^{\top}_{K}f_{K,\Sigma}(x_{0})\nabla_{K}f_{K,\Sigma}(x_{0})$
$\quad+(1-\gamma)\text{Tr}[\nabla_{\Sigma}C_{K,\Sigma}(x_{0})((R+\gamma P_{K}(B^{\top}B+D^{\top}D)))^{-1}\nabla_{\Sigma}C_{K,\Sigma}(x_{0})]$
$\leq\frac{S_{K^{*},\Sigma^{*}}}{\mu^{2}\sigma_{min}(R)}\nabla^{\top}_{K}f_{K,\Sigma}(x_{0})\nabla_{K}f_{K,\Sigma}(x_{0})+\frac{(1-\gamma)\text{Tr}[(\nabla_{\Sigma}C_{K,\Sigma}(x_{0}))^{2}]}{\sigma_{min}(R)}$
As $q_{K,\Sigma}$ is a concave function w.r.t. $\Sigma$, so $q_{K,\Sigma^{*}}-q_{K,\Sigma}\leq\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma-\Sigma^{*})$. $\Sigma\preceq I$
Now, taking the expectation w.r.t $x_{0}$ on both sides we have,
$f(K,\Sigma)-f(K^{*},\Sigma)\leq\frac{1}{\mu\sigma_{min}(R)}\nabla_{K}f^{\top}(K,\Sigma)\nabla_{K}f(K,\Sigma)+\frac{(1-\gamma)\|\nabla_{\Sigma}f(K,\Sigma)\|^{2}}{\sigma_{min}(R)}$ (40)
Lower Bound $C_{K^{*}}(x_{0})\leq C_{K^{\prime}}(x_{0})$ for any $K^{\prime}\in\mathbb{R}^{n}$, we considering when $K^{\prime}=K-(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}$
$C_{K}(x_{0})-C_{K^{*}}(x_{0})$ $\geq C_{K}(x_{0})-C_{K^{\prime}}(x_{0})$ (41)
$=-\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}A_{K}(x^{\prime}_{t},K^{\prime})\right]$ (42)
$=S_{K}[E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}]+h_{K}(\Sigma)-h_{K}(\Sigma^{\prime})$ (43)
$\geq\frac{\mathbb{E}[x_{0}^{2}]}{S_{K}^{2}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\nabla_{K}C_{K,\Sigma}^{\top}(x_{0})\nabla_{K}C_{K,\Sigma}(x_{0})$ (44)
taking the expectation w.r.t. $x_{0}$ on both sides we have,
$f(K,\Sigma)-f(K^{*},\Sigma^{*})\geq\frac{\mathbb{E}[x_{0}^{2}]}{S_{K}^{2}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\nabla f^{\top}(K,\Sigma)\nabla f(K,\Sigma)$ (45)
B.2 Proof of Lemma 3.2
Proof.
$\|\nabla_{K}f(K,\Sigma)\|$ $=\|E_{K}S_{K,\Sigma}\|$
$\leq\|E_{K}\|\frac{f(K,\Sigma)-\Omega}{Q}$
$\leq\frac{f(K,\Sigma)-\Omega}{Q}\sqrt{\lambda_{1}^{-1}(f(K,\Sigma)-f(K^{*},\Sigma^{*}))}$ (46)
Where $\Omega=Tr(\Sigma R)-\frac{\tau}{2}\left(n+log(2\pi)^{n}|\Sigma|\right)$ and the last line is by Lemma 3.1
$\|\nabla_{\Sigma}f(K,\Sigma)\|$ $=\|(1-\gamma)^{-1}\left((R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{\top}+\frac{\tau}{2}\|\Sigma^{-1}\|\right)\|$
$\leq(1-\gamma)^{-1}\left(\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2}\|\Sigma^{-1}\|\right)$
$\leq(1-\gamma)^{-1}\left(\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2\sigma_{min}(\Sigma)}\right)$
Where the last line is because $\|\Sigma^{-1}\|\leq\sigma_{max}(\Sigma^{-1})=\frac{1}{\sigma_{min}(\Sigma)}$
B.3 Proof of Lemma 3.3
Proof.
$C_{K^{\prime}}(x_{0})-C_{K}(x_{0})$ $=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[A_{K}(x^{\prime}_{t},K^{\prime})]$
$=\displaystyle S_{K^{\prime},\Sigma^{\prime}}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)E_{K}]+q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma}$
We now need to prove the smoothness of $q_{K,\Sigma}$ with respect to $\Sigma$ by showing:
$q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma}+Tr\left(\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma-\Sigma^{\prime})\right)\leq\frac{m}{2}Tr((\Sigma^{-1}\Sigma^{\prime}-I)^{2})$
Observe:
$q_{K,\Sigma}-q_{K,\Sigma^{\prime}}+Tr\left(\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma^{\prime}-\Sigma)\right)=\frac{\tau}{2(1-\gamma)}\left[log(\Sigma^{-1}\Sigma^{\prime})-Tr(\Sigma^{-1}\Sigma^{\prime}-I)\right]$ (48)
[…]
B.4 Proof of Lemma 3.4
Proof.
We will first show:
$aI\preceq\Sigma-\frac{\eta}{1-\gamma}(R-\frac{\tau}{2}\Sigma^{-1}+\gamma P_{K}(B^{\top}B+D^{\top}D))\prec I$
Let $h(y)=y+\frac{\tau}{2(1-\gamma)y}$ which is monotonic increasing for $y\in\left[\sqrt{\frac{\eta\tau}{2(1-\gamma)}},\infty\right)$ as $\sqrt{\frac{\eta\tau}{2(1-\gamma)}}\leq a\leq\frac{\sigma_{min}(R)}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}<1$. Now observe:
$\Sigma+\frac{\eta\tau}{2(1-\gamma)}\Sigma^{\prime}-\frac{\eta}{1-\gamma}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))\succeq\left(a+\frac{\eta\tau}{2(1-\gamma)a}\right)-\frac{\eta}{(1-\gamma)}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))$
$\succeq\left(a+\frac{\eta}{1-\gamma}\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|\right)I-\frac{\eta}{1-\gamma}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))\succeq aI$
Now,
$\Sigma+\frac{\eta\tau}{2(1-\gamma)}\Sigma^{-1}-\frac{\eta}{1-\gamma}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))\preceq\left(1+\frac{\eta\tau}{2(1-\gamma)}\right)I-\frac{\eta}{1-\gamma}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))$
$\preceq\left(1+\frac{\eta}{1-\gamma}\sigma_{min}(R)\right)I-\frac{\eta}{1-\gamma}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))\preceq I$
And so using these facts we can write,
$aI\preceq\Sigma-\frac{\eta}{1-\gamma}(R-\frac{\tau}{2}\Sigma^{-1}+\gamma P_{K}(B^{\top}B+D^{\top}D))\prec I$
Next I will show $aI\preceq\Sigma^{\prime}\preceq I$. Observe that:
$aI-\Sigma\preceq-\frac{\eta}{1-\gamma}(R+\gamma P_{K}(B^{\top}B+D^{\top}D)-\frac{\tau}{2}\Sigma^{-1})\preceq I-\Sigma$
Now multiply both sides by $\Sigma$ and add $\Sigma$ which yields:
$a\Sigma^{2}-\Sigma^{3}+\Sigma\preceq\Sigma-\frac{\eta}{1-\gamma}\Sigma\left(R+\gamma P_{K}(B^{\top}B+D^{\top}D)-\frac{\tau}{2}\Sigma^{-1}\right)\Sigma\preceq\Sigma^{2}-\Sigma^{3}+\Sigma$
$a<\sigma_{min}(\Sigma)$ so we have $aI-\Sigma\preceq 0$. $aI-\Sigma\preceq(aI-\Sigma)\Sigma^{2}$ as $\Sigma\preceq I$. $a\Sigma^{2}-\Sigma^{3}+\Sigma\succeq aI$.
As $I-\Sigma\succeq 0$, we have $I-\Sigma\succeq(I-\Sigma)\Sigma^{2}$, so $\Sigma^{2}-\Sigma^{3}+\Sigma\preceq I$ The proof is completed ∎
B.5 Proof of Lemma 3.5
Proof.
$f(K^{\prime},\Sigma^{\prime})-f(K,\Sigma)$ $=S_{K^{\prime}}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
$\quad+q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma}$
Using RPG and $\eta_{1}\leq\frac{1}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ we have
$S_{K^{\prime}}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
$\quad\leq S_{K^{\prime}}[\eta_{1}^{2}E_{K}^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))E_{K}-2\eta_{1}E_{K}^{\top}E_{K}]$
$\quad\leq-\eta_{1}S_{K^{\prime}}E_{K}^{\top}E_{K}$
$\quad\leq-\eta_{1}\mu\frac{\sigma_{min}{(R)}}{S_{K^{*},\Sigma^{*}}}E^{\top}_{K}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}$
From lemma 3.3, we have
$q_{K,\Sigma^{\prime}}-q_{K,\Sigma}$ $\leq\frac{\text{Tr}\left(((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}\Sigma^{-1})(\Sigma^{\prime}-\Sigma)\right)}{(1-\gamma)}+\frac{\tau m}{2(1-\gamma)}Tr((\Sigma^{-1}\Sigma^{\prime}-I)^{2})$
$=\frac{-\eta_{2}}{{(1-\gamma)}}{\text{Tr}[\left(((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}\Sigma^{-1})\Sigma\right)^{2}]}$
$\quad+\frac{\eta_{2}^{2}\tau m}{2(1-\gamma)^{3}}\text{Tr}(((R+\gamma P_{K}(B^{\top}B+D^{\top}D))\Sigma-\frac{\tau}{2}I)^{2})$
$\leq-\frac{\eta_{2}}{2(1-\gamma)^{2}}\text{Tr}[(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{2}]$
$\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}\leq\frac{2(1-\gamma)}{\tau}\left(\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}\right)\leq\frac{2(1-\gamma)}{\tau}\leq\frac{(1-\gamma)}{\tau m}$
From lemma 3.1, we have
$q_{K,\Sigma^{*}}-q_{K,\Sigma}$ $\leq Tr[\nabla_{\Sigma}C_{K,\Sigma}(x_{0})((R+\gamma P_{K}(B^{\top}B+D^{\top}D)))^{-1}((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}I)]$
$\leq\frac{1}{(1-\gamma)\sigma_{min}(R)}Tr[((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}I)^{2}\Sigma^{-1}]$
$\leq\frac{1}{(1-\gamma)a\sigma_{min}(R)}Tr[((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}I)^{2}]$
Combining the above we have
$q_{K,\Sigma^{\prime}}-q_{K,\Sigma}\leq\frac{\eta_{2}a\sigma_{min}(R)}{2(1-\gamma)}(q_{K,\Sigma}-q_{K,\Sigma^{*}})$
Finally, with $\phi=min\{\eta_{1}\mu\frac{\sigma_{min}{(R)}}{S_{K^{*},\Sigma^{*}}},\frac{\eta_{2}a\sigma_{min}(R)}{2(1-\gamma)}\}$, we have
$f(K^{\prime},\Sigma^{\prime})-f(K,\Sigma)$
$\quad\leq-\eta_{1}\mu\frac{\sigma_{min}{(R)}}{S_{K^{*},\Sigma^{*}}}E^{\top}_{K}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}+\frac{\eta_{2}\sigma_{min}(R)}{2(1-\gamma)}(q_{K,\Sigma}-q_{K,\Sigma^{*}})$
$\quad\leq-\phi\left({S_{K^{*},\Sigma^{*}}}E^{\top}_{K}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))^{-1}E_{K}+(q_{K,\Sigma^{*}}-q_{K,\Sigma})\right)$
$\quad\leq-\phi(f(K,\Sigma)-f(K^{*},\Sigma^{*}))$
$f(K^{\prime},\Sigma^{\prime})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{*},\Sigma^{*}))$
B.6 Proof of Lemma 3.6
Proof.
$q_{K,\Sigma}$ $=\frac{\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|)))}{1-\gamma}$
$\geq\frac{1}{1-\gamma}[\sigma_{min}(R)\text{Tr}(\Sigma)-\frac{\tau}{2}(k+klog(2\pi)+\log|\Sigma|)]$
$\geq\frac{1}{1-\gamma}\left[\frac{\tau k}{2}-\frac{\tau}{2}(k+log(2\pi))-\frac{\tau K}{2}log\left(\frac{\tau}{2\sigma_{min}(R)}\right)\right].$
As $\frac{\tau k}{2}-\frac{\tau}{2}(k+log(2\pi))-\frac{\tau k}{2}log(\frac{\tau}{2\sigma_{min}(R)})$ is a convex function w.r.t. $\Sigma$ with minimizer $\frac{\tau}{2\sigma_{min}(R)}I$, so we have,
$q_{K,\Sigma}\geq\frac{\tau k}{2(1-\gamma)}log\left(\frac{\sigma_{min}(R)}{\pi\tau}\right)$
Appendix C Proofs in Section 4
C.1 Proof of Lemma 4.1
Proof.
$f(K,\Sigma)$ $=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(Qx_{t}^{2}+u_{t}^{T}Ru_{t})\right]-\frac{\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}$
$=(Q+K^{\top}RK)S_{K}+\frac{Tr(\Sigma R)-\frac{\tau}{2}\left(n+log(2\pi)^{n}|\Sigma|\right)}{1-\gamma}.$
Then we have
$S_{K}$ $=\frac{f(K,\Sigma)-(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}\left(n+log(2\pi)^{n}|\Sigma|\right)\right]}{Q+K^{\top}RK}$
$\leq\frac{f(K,\Sigma)-(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}\left(n+log(2\pi)^{n}|\Sigma|\right)\right]}{Q}$
$\mathbb{E}[x^{2}_{t+1}]$ $=V_{K}\mathbb{E}[x_{t}^{2}]+Tr(\Sigma(B^{\top}B+D^{\top}D))$
$=V_{K}(V_{K}\mathbb{E}[x_{t-1}^{2}]+Tr(\Sigma(B^{\top}B+D^{\top}D)))+Tr(\Sigma(B^{\top}B+D^{\top}D)))$
$=V_{K}^{2}\mathbb{E}[x_{t-1}^{2}]+(V_{K}+1)Tr(\Sigma(B^{\top}B+D^{\top}D))$
$=V_{K}^{t+1}\mu+Tr(\Sigma(B^{\top}B+D^{\top}D))\sum_{i=0}^{t}V_{K}^{i}$
So observe:
$S_{K}$ $=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[x_{t}^{2}]$
$=\sum_{t=0}^{\infty}\gamma^{t}\left(V_{K}^{t}\mathbb{E}[x_{0}^{2}]+Tr(\Sigma(B^{\top}B+D^{\top}D))\frac{1-V_{K}^{t}}{1-V_{K}}\right)$
$=\mu\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{\infty}\gamma^{t}(1-V_{K}^{t})$
$=\mu\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{1}{1-\gamma}-\frac{1}{1-\gamma V_{K}}\right]$
C.2 Proof of Lemma 4.2
Proof.
Note that
$f(K,\Sigma)$ $=\mathbb{E}[P_{K}x_{0}^{2}+q_{K,\Sigma}]$
$=(Q+K^{\top}RK)S_{K}+(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]$
$S_{K}$ $=\mathbb{E}[x_{0}^{2}]\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{\infty}\gamma^{t}(1-V_{K}^{t})$ (49)
So
$f^{(l)}(K,\Sigma)$ $=(Q+K^{\top}RK)S^{(l)}_{K}+\frac{1-\gamma^{l}}{1-\gamma}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]$
$S^{(l)}_{K,\Sigma}$ $=\mu\sum_{t=0}^{l-1}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{l-1}\gamma^{t}(1-V_{K}^{t})$ (50)
Next note that
$S_{K}-S_{K}^{(l)}$ $=\mu\sum_{t=l}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=l}^{\infty}\gamma^{t}(1-V_{K}^{t})$
$=(\mu-\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}})\frac{(\gamma V_{K})^{l}}{1-\gamma V_{K}}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\frac{\gamma^{l}}{1-\gamma}$
$\leq(\mu-\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}})\frac{\gamma^{l}}{1-\gamma V_{K}}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\frac{\gamma^{l}}{1-\gamma}$
$=\gamma^{l}S_{K,\Sigma}$
and
$f(K,\Sigma)-f^{(l)}(K,\Sigma)$ $=(Q+K^{\top}RK)\sum_{t=l}^{\infty}\mathbb{E}\gamma^{t}x_{t}^{2}+(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]\gamma^{l}$
$=(Q+K^{\top}RK)(S_{K}-S_{K}^{(l)})+(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]\gamma^{l}$
$\leq(Q+K^{\top}RK)\gamma^{l}S_{K,\Sigma}+(1-\gamma)^{-1}\left[Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)\right]\gamma^{l}$ (51)
$=\gamma^{l}\left[(Q+K^{\top}RK)S_{K,\Sigma}+\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\right]$ (52)
taking $l$ in the above inequality completes the proof. ∎
C.3 Two useful Lemmas
Lemma C.1.
If $\frac{|\gamma V_{K}-\gamma V_{K^{\prime}}|}{1-\gamma V_{K}}\leq\frac{1}{2}$, then
$|(1-\gamma V_{K})^{-1}-(1-\gamma V_{K^{\prime}})^{-1}|$ $\leq 2(1-\gamma V_{K})^{-2}|\gamma V_{K^{\prime}}-\gamma V_{K}|$
Proof.
We have
$(1-(1-\gamma V_{K})^{-1}(\gamma V_{K^{\prime}}-\gamma V_{K}))^{-1}\leq(1-(1-\gamma V_{K})^{-1}|\gamma V_{K^{\prime}}-\gamma V_{K}|)^{-1}\leq 2$
as $(1-\gamma V_{K})^{-1}|\gamma V_{K}-\gamma V_{K^{\prime}}|\leq\frac{1}{2}$.
$(1-\gamma V_{K})^{-1}-(1-\gamma V_{K^{\prime}})^{-1}$ $=(1-\gamma V_{K})^{-1}-[(1-\gamma V_{K})-(\gamma V_{K^{\prime}}-\gamma V_{K})]^{-1}$
$=(1-\gamma V_{K})^{-1}[1-(1-(1-\gamma V_{K})^{-1}(\gamma V_{K^{\prime}}-\gamma V_{K}))^{-1})]$
$|1-(1-(1-\gamma V_{K})^{-1}(\gamma V_{K^{\prime}}-\gamma V_{K}))^{-1}|$ $=(1-\gamma V_{K})^{-1}|(\gamma V_{K^{\prime}}-\gamma V_{K})(1-(1-\gamma V_{K})^{-1}(\gamma V_{K^{\prime}}-\gamma V_{K}))^{-1}|$
$\leq 2(1-\gamma V_{K})^{-1}|\gamma V_{K^{\prime}}-\gamma V_{K}|$
So we have $(1-\gamma V_{K})^{-1}-(1-\gamma V_{K^{\prime}})^{-1}\leq 2(1-\gamma V_{K})^{-2}|\gamma V_{K^{\prime}}-\gamma V_{K}|$
Lemma C.2.
If $\|K-K^{\prime}\|\leq h_{\Sigma}$,then
$|\frac{1}{1-\gamma V_{K^{\prime}}}-\frac{1}{1-\gamma V_{K}}|\leq g_{\Sigma}\|K^{\prime}-K\|,$
where $h_{\Sigma}$ is defined in Lemma 4.3
Proof.
$|\gamma V_{K}-\gamma V_{K^{\prime}}|$ $=(A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK)$
$\quad-(A^{2}+C^{2}+K^{\prime\top}(B^{\top}B+D^{\top}D)K^{\prime}-2ABK^{\prime})$
$=K^{\top}(B^{\top}B+D^{\top}D)K-2AB(K-K^{\prime})$
$\quad-(K+K^{\prime}-K)^{\top}(B^{\top}B+D^{\top}D)(K+K^{\prime}-K)$
$=-(K-K^{\prime})^{\top}(B^{\top}B+D^{\top}D)(K-K^{\prime})-2AB(K-K^{\prime})$
$\quad+2K^{\top}(B^{\top}B+D^{\top}D)(K-K^{\prime})$
$\leq 2\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|\|K-K^{\prime}\|$
$\quad+\|(B^{\top}B+D^{\top}D)\|\|K-K^{\prime}\|^{2}$
$=\|K-K^{\prime}\|(2\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|+\|B^{\top}B+D^{\top}D\|\|K-K^{\prime}\|)$
$=\|B^{\top}B+D^{\top}D\|\left(\|K-K^{\prime}\|+\frac{\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}{\|B^{\top}B+D^{\top}D\|}\right)^{2}$
$\quad-\frac{\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|^{2}}{\|B^{\top}B+D^{\top}D\|}$
If we set
$\|K-K^{\prime}\|$ $\leq\sqrt{\frac{1}{2}\frac{(1-\gamma V_{K})^{2}}{\|B^{\top}B+D^{\top}D\|}+\left(\frac{\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}{\|B^{\top}B+D^{\top}D\|}\right)^{2}}-\frac{\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}{\|B^{\top}B+D^{\top}D\|}$
$=\frac{\frac{1}{2}\frac{(1-\gamma V_{K})^{2}}{\|B^{\top}B+D^{\top}D\|}}{\sqrt{\frac{1}{2}\frac{(1-\gamma V_{K})^{2}}{\|B^{\top}B+D^{\top}D\|}+\left(\frac{\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}{\|B^{\top}B+D^{\top}D\|}\right)^{2}}+\frac{\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}{\|B^{\top}B+D^{\top}D\|}}$ (53)
$=\frac{\frac{1}{2}{(1-\gamma V_{K})^{2}}}{\sqrt{\frac{1}{2}{(1-\gamma V_{K})^{2}\|B^{\top}B+D^{\top}D\|}+\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|^{2}}+\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}$
$=\frac{1}{2}\frac{(1-\gamma V_{K})^{2}}{\sqrt{\frac{1}{2}(1-\gamma V_{K})^{2}\|B^{\top}B+D^{\top}D\|+\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|^{2}}+\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|}$
$:=h_{\Sigma}$ (54)
then we have
$\frac{1}{1-\gamma V_{K}}|\gamma V_{K}-\gamma V_{K^{\prime}}|$ $\leq\frac{1}{1-\gamma V_{K}}\frac{(1-\gamma V_{K})^{2}}{2}$
$\leq\frac{1-\gamma V_{K}}{2}$
$\leq\frac{1}{2}$ (55)
which satisfies the condition in lemma C.1. Apply lemma C.1 we have,
$|\frac{1}{1-\gamma V_{K}}-\frac{1}{1-\gamma V_{K^{\prime}}}|$ $\leq 2\left(\frac{1}{1-\gamma V_{K}}\right)^{2}|\gamma V_{K}-\gamma V_{K^{\prime}}|$
$\leq 2\left(\frac{1}{1-\gamma V_{K}}\right)^{2}(2\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|+\|B^{\top}B+D^{\top}D\|\|K-K^{\prime}\|)\|K-K^{\prime}\|$
$\leq 2\left(\frac{1}{1-\gamma V_{K}}\right)^{2}(2\|K^{\top}(B^{\top}B+D^{\top}D)-AB\|+\|B^{\top}B+D^{\top}D\|h_{\Sigma})\|K-K^{\prime}\|$
$:=g_{\Sigma}\|K-K^{\prime}\|$
C.4 Proof of Lemma 4.3
Proof.
We have $\|\Sigma^{\prime}\|\leq 2\|\Sigma\|$ as $\|\Sigma^{\prime}\|-\|\Sigma\|\leq\|\Sigma^{\prime}-\Sigma\|\leq\|\Sigma\|$
$S_{K,\Sigma}$ $=\mu\sum_{t=0}^{\infty}(\gamma V_{K})^{t}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\sum_{t=0}^{\infty}\gamma^{t}(1-V_{K}^{t})$
$=\frac{\mu}{1-\gamma V_{K}}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{1}{1-\gamma}-\frac{1}{1-\gamma V_{K}}\right]$
$=\frac{\mu}{1-\gamma V_{K}}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{(1-\gamma V_{K})-(1-\gamma)}{(1-\gamma)(1-\gamma V_{K})}\right]$
$=\frac{\mu}{1-\gamma V_{K}}+\frac{Tr(\Sigma(B^{\top}B+D^{\top}D))}{1-V_{K}}\left[\frac{\gamma(1-V_{K})}{(1-\gamma)(1-\gamma V_{K})}\right]$
$=\frac{(1-\gamma)\mu+\gamma Tr(\Sigma(B^{\top}B+D^{\top}D))}{(1-\gamma)(1-\gamma V_{K})}$
$:=\Delta_{\Sigma}(1-\gamma V_{K})^{-1}$
$|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|$ $\leq|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma^{\prime}}|+|S_{K,\Sigma^{\prime}}-S_{K,\Sigma}|$
$\leq\Delta_{\Sigma^{\prime}}|\frac{1}{1-\gamma V_{K^{\prime}}}-\frac{1}{1-\gamma V_{K}}|+\frac{\gamma Tr((\Sigma^{\prime}-\Sigma)(B^{\top}B+D^{\top}D))}{(1-\gamma)(1-\gamma V_{K})}$
$\leq\Delta_{\Sigma^{\prime}}g_{\Sigma}\|K^{\prime}-K\|+\frac{\gamma Tr((B^{\top}B+D^{\top}D))}{(1-\gamma)(1-\gamma V_{K})}\|\Sigma^{\prime}-\Sigma\|$
$\leq 2\Delta_{\Sigma}g_{\Sigma}\|K^{\prime}-K\|+\frac{\gamma Tr((B^{\top}B+D^{\top}D))}{(1-\gamma)(1-\gamma V_{K})}\|\Sigma^{\prime}-\Sigma\|$
$:=h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma-\Sigma^{\prime}\|$
C.5 Proof of Lemma 4.4
Proof.
From $\|K^{\prime}\|-\|K\|\leq\|K-K^{\prime}\|\leq|K|$ we have $K^{\prime}\leq 2\|K\|$ and $K^{\prime}RK^{\prime}\leq\|R\|\|K^{\prime}\|^{2}\leq 4\|R\|\|K\|^{2}$
$|P_{K}-P_{K^{\prime}}|$ $=\bigg|\frac{Q+K^{\top}RK}{1-\gamma V_{K}}-\frac{Q+K^{\prime\top}RK^{\prime}}{1-\gamma V_{K^{\prime}}}\bigg|$
$\leq\bigg|\frac{Q+K^{\top}RK}{1-\gamma V_{K}}-\frac{Q+K^{\prime\top}RK^{\prime}}{1-\gamma V_{K}}\bigg|+\bigg|\frac{Q+K^{\prime\top}RK^{\prime}}{1-\gamma V_{K}}-\frac{Q+K^{\prime\top}RK^{\prime}}{1-\gamma V_{K^{\prime}}}\bigg|$
$=\frac{|{K^{\prime}}^{\top}RK^{\prime}-K^{\top}RK|}{1-\gamma V_{K}}+(Q+{K^{\prime}}^{\top}RK^{\prime})\bigg|\frac{1}{1-\gamma V_{K}}-\frac{1}{1-\gamma V_{K^{\prime}}}\bigg|$
$\leq\frac{|{K^{\prime}}^{\top}RK^{\prime}-K^{\top}RK|}{1-\gamma V_{K}}+(Q+4\|R\|\|K\|^{2})g_{\Sigma}\|K-K^{\prime}\|$
$=\frac{|({K^{\prime}}-K)^{\top}R(K^{\prime}-K)-2K^{\top}R(K^{\prime}-K)|}{1-\gamma V_{K}}+(Q+4\|R\|\|K\|^{2})g_{\Sigma}\|K-K^{\prime}\|$
$\leq\frac{3\|K\|\|R\|}{1-\gamma V_{K}}\|K^{\prime}-K\|+(Q+4\|R\|\|K\|^{2})g_{\Sigma}\|K-K^{\prime}\|$
$:=h_{5}\|K^{\prime}-K\|$
C.6 Proof of Lemma 4.5
Proof.
We first consider $\nabla_{K}f(K,\Sigma)$ perturbation
$\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)$ $=E_{K^{\prime}}S_{K^{\prime},\Sigma^{\prime}}-E_{K}S_{K,\Sigma}$
$=(E_{K^{\prime}}-E_{K})S_{K^{\prime},\Sigma^{\prime}}-(S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma})E_{K}$
$\leq(E_{K^{\prime}}-E_{K})S_{K^{\prime},\Sigma^{\prime}}+(S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma})E_{K}$
By triangle equality we have
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|$ $\leq\|E_{K^{\prime}}-E_{K}\||S_{K^{\prime},\Sigma^{\prime}}|+|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|E_{K}$
By definition of $E_{K}$, we have
$\frac{1}{2}\|E_{K^{\prime}}-E_{K}\|$ $=\|R\|\|K^{\prime}-K\|+\gamma P_{K^{\prime}}[(B^{\top}B+D^{\top}D)K^{\prime}]-\gamma P_{K}[(B^{\top}B+D^{\top}D)K]$
$\quad+\gamma|P_{K^{\prime}}-P_{K}|A\|B\|$
$\leq\|R\|\|K^{\prime}-K\|+\gamma A\cdot h_{5}\|B\|\|K^{\prime}-K\|+\gamma P_{K^{\prime}}(B^{\top}B+D^{\top}D)K^{\prime}$
$\quad-\gamma P_{K}(B^{\top}B+D^{\top}D)(K-K^{\prime}+K^{\prime})$
$\leq\|R\|\|K^{\prime}-K\|+\gamma A\cdot h_{5}\|B\|\|K^{\prime}-K\|+\gamma P_{K}(B^{\top}B+D^{\top}D)\|K^{\prime}-K\|$
$\quad+\gamma|P_{K^{\prime}}-P_{K}|\|B^{\top}B+D^{\top}D\|\|K^{\prime}\|$
$\leq\|R\|\|K^{\prime}-K\|+\gamma A\cdot h_{5}\|B\|\|K^{\prime}-K\|+\gamma P_{K}(B^{\top}B+D^{\top}D)\|K^{\prime}-K\|$
$\quad+2\gamma\cdot h_{5}\|B^{\top}B+D^{\top}D\|\|K\|\|K^{\prime}-K\|$
$=\left[\|R\|+\gamma A\cdot h_{5}\|B\|+\gamma P_{K}(B^{\top}B+D^{\top}D)++2\gamma\cdot h_{5}\|B^{\top}B+D^{\top}D\|\|K\|\right]\|K^{\prime}-K\|$
$\|E_{K^{\prime}}-E_{K}\|$ $\leq 2\left[\|R\|+\gamma A\cdot h_{5}\|B\|+\gamma P_{K}(B^{\top}B+D^{\top}D)++2\gamma\cdot h_{5}\|B^{\top}B+D^{\top}D\|\|K\|\right]\|K^{\prime}-K\|$
$:=h_{E}\|K^{\prime}-K\|$
Using $S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}\leq|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|$ and lemma 4.4 we have
$S_{K^{\prime},\Sigma^{\prime}}$ $\leq S_{K,\Sigma}+h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma^{\prime}-\Sigma\|$
$\leq S_{K,\Sigma}+h_{K}h_{\Sigma}+h_{2}\|\Sigma\|,$
and
$\|E_{K^{\prime}}-E_{K}\|S_{K^{\prime},\Sigma^{\prime}}\leq(S_{K,\Sigma}+h_{K}h_{\Sigma}+h_{2}\|\Sigma\|)h_{E}\|K^{\prime}-K\|$
if $\|K^{\prime}-K\|\leq h_{\Sigma}$ and $\|\Sigma-\Sigma^{\prime}\|\leq\|\Sigma\|$.
From Lemma 3.1, we have
$\|E_{K}\|\leq\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}$
so for the second term we have:
$|S_{K^{\prime},\Sigma^{\prime}}-S_{K,\Sigma}|\|E_{K}\|\leq(h_{K}\|K^{\prime}-K\|+h_{2}\|\Sigma^{\prime}-\Sigma\|)\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}$ (56)
by lemma 4.3. Then we have
$\|\nabla_{K}f(K^{\prime},\Sigma^{\prime})-\nabla_{K}f(K,\Sigma)\|$ $\leq\left(h_{K}\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}+(S_{K,\Sigma}+h_{K}h_{\Sigma}+h_{2}\|\Sigma\|)h_{E}\right)\|K^{\prime}-K\|$
$+\left(h_{2}\sqrt{\lambda_{1}^{-1}|f(K,\Sigma)-f(K^{*},\Sigma^{*})|}\right)\|\Sigma^{\prime}-\Sigma\|$
Now consider the perturbation of $\nabla_{\Sigma}f(K,\Sigma)$
Apply theorem 35 in [9] we have
$\|\Sigma^{\prime-1}-\Sigma^{-1}\|$ $\leq\frac{2}{\sigma_{min}(\Sigma)}\|\Sigma-\Sigma^{\prime}\|$
if $\|\Sigma-\Sigma^{\prime}\|\leq\frac{\sigma_{min}(\Sigma)}{2}$
$\nabla_{\Sigma}f(K^{\prime},\Sigma^{\prime})-\nabla_{\Sigma}f(K,\Sigma)$ $=\frac{\left((R+\gamma P_{K^{\prime}}(B^{\top}B+D^{\top}D))^{\top}-\frac{\tau}{2}\Sigma^{\prime-1}\right)-\left((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}\Sigma^{-1}\right)}{1-\gamma}$
$=\frac{\gamma\left(P_{K^{\prime}}-P_{K}\right)(B^{\top}B+D^{\top}D)-\frac{\tau}{2}(\Sigma^{\prime-1}-\Sigma^{-1})}{(1-\gamma)}$
So we have,
$\|\nabla_{\Sigma}f(K^{\prime},\Sigma^{\prime})-\nabla_{\Sigma}f(K,\Sigma)\|$ $\leq\frac{\gamma(B^{\top}B+D^{\top}D)}{(1-\gamma)}|P_{K^{\prime}}-P_{K}|+\frac{\tau}{2(1-\gamma)}\|\Sigma^{\prime-1}-\Sigma^{-1}\|$
$\leq\frac{\gamma(B^{\top}B+D^{\top}D)}{(1-\gamma)}h_{5}\|K^{\prime}-K\|+\frac{\tau\sigma_{min}(\Sigma)}{4(1-\gamma)}\|\Sigma-\Sigma^{\prime}\|$
$\leq\frac{\gamma(B^{\top}B+D^{\top}D)}{(1-\gamma)}h_{5}\|K^{\prime}-K\|+\frac{\tau\sigma_{min}(\Sigma)}{4(1-\gamma)}\|\Sigma^{\prime}-\Sigma\|_{F}$
$:=h_{8}\|K^{\prime}-K\|+h_{9}\|\Sigma^{\prime}-\Sigma\|_{F}$
The proof is completed ∎
C.7 Proof of Lemma 4.6
Proof.
The proof is provided in Lemma 2.1 of [24] with slightly change of notation. ∎
C.8 Proof of Lemma 4.7
Proof.
We first show that finitely many finite-horizon rollouts, defined as
$\widetilde{\nabla}_{K}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r^{2}}f(K+U_{i},\Sigma)U_{i}.$
If $r_{1}$ satisfies the conditions in the lemma, then $\|\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)\|_{F}<\frac{\epsilon}{3}$ with high probability(at least $1-\sqrt{\kappa_{1}}$). We break $\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)$ into the following two parts,
$\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)=(\nabla_{K}f_{r_{1}}(K,\Sigma)-\nabla_{K}f(K,\Sigma))+(\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma))$
For the first term, from Lemma 4.5 we have $\|\nabla_{K}f(K+U,\Sigma)-\nabla_{K}f(K,\Sigma)\|\leq\frac{\epsilon}{6}$ if we set $r_{1}\leq\frac{\epsilon}{6h_{6}}$. Since $\nabla_{K}f_{r_{1}}(K,\Sigma)$ is the expectation of $\nabla f(K+U,\Sigma)$, we have $\|\nabla_{K}f_{r_{1}}(K,\Sigma)-\nabla_{K}f(K,\Sigma)\|_{F}\leq\frac{\epsilon}{6}$.
For the second term, $\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)$. We want to invoke the Vector Bernstein Inequality to show that with high probability $\hat{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)||<\frac{\epsilon}{2}$. Consider the sample $i^{th}$ of a single path $K+U_{i}$ and observe that $\|\frac{n}{r_{1}^{2}}f(K+U_{i})U_{i}\|\leq\frac{2nf(K,\Sigma)}{r_{1}}$ if we assume $f(K+U_{i},\Sigma)\leq 2f(K,\Sigma)$. Also,
$\|\nabla_{K}f_{r_{1}}(K,\Sigma)\|$ $\leq\|\nabla_{K}f_{r_{1}}(K,\Sigma)-\nabla_{K}f(K,\Sigma)\|_{F}+\nabla_{K}f(K,\Sigma)$
$\leq\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
So we have
$\|\frac{n}{r_{1}^{2}}f_{r_{1}}(K+U_{i},\Sigma)U_{i}-\nabla_{K}f_{r_{1}}(K,\Sigma)\|\leq R_{1}:=\frac{2n}{r_{1}}f(K,\Sigma)+\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
and
$\|\mathbb{E}\bigg[\big(\frac{n}{r_{1}^{2}}f_{r_{1}}(K+U_{i},\Sigma)U_{i}\big)^{\top}\frac{n}{r_{1}^{2}}f_{r_{1}}(K+U_{i},\Sigma)U_{i}\bigg]-\nabla^{\top}_{K}f_{r_{1}}(K,\Sigma)\nabla_{K}f_{r_{1}}(K,\Sigma)\|$
$\quad\leq\max_{U_{i}}\|\frac{n}{r_{1}^{2}}f_{r_{1}}(K+U_{i},\Sigma)U_{i}\|_{F}^{2}+\|\nabla_{K}f_{r_{1}}(K,\Sigma)\|^{2}_{F}$
$\quad\leq\sigma_{1}:=\big(\frac{2n}{r_{1}}f(K,\Sigma)\big)^{2}+\big(\frac{\epsilon}{6}+\overline{\|\nabla_{K}f(K,\Sigma)\|}\big)^{2}$
Next, note that $\mathbb{E}{[\frac{n}{r_{1}^{2}}f(K+U_{i},\Sigma)U_{i}]}=\mathbb{E}{[\hat{\nabla}_{K}]}=\nabla_{K}f_{r_{1}}(K,\Sigma)$, apply Vector Bernstein Inequality we have if
$M\geq\frac{2n}{(\epsilon_{1}/6)^{2}}(\sigma_{\nabla}+\frac{R_{\nabla}}{3\sqrt{n}})\log\big(\frac{n+1}{\sqrt{\kappa_{1}}}\big)$
then
$\mathbb{P}\big[\|\widetilde{\nabla}_{K}-\nabla_{K}f_{r_{1}}(K,\Sigma)\|_{F}\leq\frac{\epsilon}{6}\big]\geq 1-\sqrt{\kappa_{1}}.$
In the above, we have demonstrated that $\|\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)\|_{F}\leq\frac{\epsilon}{3}$ with high probabiliy. Now we attempt to prove that $\widehat{\nabla}_{K}$ is $\epsilon$ close to $\nabla_{K}f(K,\Sigma)$ with high probability $1-\kappa_{1}$, under the conditions given in the lemma. Define
$\nabla_{K}^{\prime}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r^{2}_{1}}f^{(l)}(K+U_{i},\Sigma)U_{i}.$
We break $\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)$ into following three parts,
$\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)=(\widehat{\nabla}_{K}-\nabla_{K}^{\prime})+(\nabla_{K}^{\prime}-\widetilde{\nabla}_{K})+(\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma))$
For third term, based on the previous proof, we know that $\|\widetilde{\nabla}_{K}-\nabla_{K}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with probability at least $1-\kappa_{1}$ under the conditions given in the lemma.
For the second term, using the lemma 4.2 we have
[…]
$\leq\frac{1}{M}\cdot\frac{n}{r_{1}^{2}}\left(\sum_{i=1}^{M}\gamma^{l}\left[(|Q|\cdot|S_{K+U_{i},\Sigma}|+\|K+U_{i}\|^{2}\|R\|\cdot|S_{K+U_{i},\Sigma}|+\Bigg|\frac{Tr(\Sigma R)-\frac{\tau}{2}(n+log(2\pi)^{n}|\Sigma|)}{1-\gamma}\Bigg|\right]r_{1}\right)$
$\leq\frac{1}{M}\cdot\frac{n}{r_{1}}\left(M\gamma^{l}\left[|2f(K,\Sigma)-\psi|+2\|K\|^{2}\|R\|\cdot\Bigg|\frac{2f(K,\Sigma)-\psi}{Q}\Bigg|+\Bigg|\frac{\psi}{1-\gamma}\Bigg|\right]\right)$
$\leq\frac{n}{r_{1}}\left(\gamma^{l}\left[|2f(K,\Sigma)|+|\psi|+2\|K\|^{2}\|R\|\cdot\frac{|2f(K,\Sigma)|+|\psi|}{|Q|}\Bigg|+\Bigg|\frac{\psi}{1-\gamma}\Bigg|\right]\right)$
$\leq\frac{\epsilon}{3},$
if $l\geq\log(\gamma)^{-1}\left[\log\left(\frac{r_{1}}{n}\cdot\frac{\epsilon}{3}\right)-\log\left(2|f(K,\Sigma)|\left(2\|K\|^{2}\|R\|+\frac{1}{|Q|}\right)+|\psi|\left(1+\frac{1}{|Q|}+\frac{1}{1-\gamma}\right)\right)\right]$, where $\psi=Tr(\Sigma R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma|)$.
For the first term, note that $|x_{0}^{i}|\leq L$ and let $\Gamma>1$ such that
$\sum_{j=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})\leq\Gamma\mathbb{E}\left[\sum_{i=0}^{l-1}Qx_{i}^{2}+u_{t}^{\top}Ru_{t}\right]$
Thus, $\widehat{\nabla}_{K}-\nabla^{\prime}_{K}$ is the sum of random bounded vectors. Now observe:
$\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})$ $\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{l-1}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\right]$
$\leq\Gamma\mathbb{E}\left[\sum_{t=0}^{\infty}Qx_{t}^{2}+u_{t}^{\top}Ru_{t}+\tau\log\pi(u_{t}|x_{t})\right]$
$\leq\Gamma L^{2}f(K+U_{i},\Sigma)$
$\leq 2\Gamma L^{2}f(K,\Sigma)$ (57)
And since
$\|\nabla_{K}^{\prime}\|\leq\frac{\epsilon}{3}+\|\widetilde{\nabla}_{K}\|\leq\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
we have
$\|\big[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\big]U_{i}-\nabla_{K}^{\prime}\|_{F}$ $\leq 2\Gamma L^{2}f(K,\Sigma)\|U_{i}\|_{F}+\|\nabla_{K}^{\prime}\|_{F}$
$\leq R_{2}$
$:=2\Gamma L^{2}f(K,\Sigma)r_{1}+\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|}$
Define $Z_{i}:=\big[\sum_{t=0}^{l-1}Q(x_{t}^{i})^{2}+(u_{t}^{i})^{\top}R(u_{t}^{i})+\tau\log\pi(u_{t}^{i}|x_{t}^{i})\big]U_{i}$
$\|\mathbb{E}[Z_{i}^{\top}Z_{i}]-(\nabla_{K}^{\prime})^{\top}\nabla_{K}^{\prime}\big]\|$ $\leq\max_{U_{i}}\|Z_{i}\|_{F}^{2}+\|\nabla_{K}^{\prime}\|_{F}^{2}$
$\leq\sigma_{2}$
$:=(2\Gamma L^{2}f(K,\Sigma)r_{1})^{2}+(\frac{\epsilon}{2}+\overline{\|\nabla_{K}f(K,\Sigma)\|})^{2}$
the norm of each sample is bounded, assuming the variance is bounded, and since $\mathbb{E}[\widehat{\nabla}_{K}]=\nabla^{\prime}_{K}$. Then by the Vector Bernstein Inequality we have if
$M\geq\frac{2n}{(\epsilon/3)^{2}}(\sigma^{2}_{2}+\frac{R_{2}\epsilon}{3\sqrt{n}})\log(\frac{n+1}{\sqrt{\kappa_{1}}})$
then
$\mathbb{P}\left(\|\widehat{\nabla}_{K}-\nabla_{K}^{\prime}\|\leq\frac{\epsilon}{3}\right)\leq 1-\sqrt{\kappa_{1}}$
Combining the above we have
$\mathbb{P}\left(\|\widehat{\nabla}_{K}-\nabla_{K}f(K,\Sigma)\|\leq{\epsilon}\right)\leq 1-\kappa_{1}$
C.9 Proof of Lemma 4.8
Proof.
We first show that finitely many finite-horizon rollouts, defined as
$\widetilde{\nabla}_{\Sigma}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r^{2}}f(K,\Sigma+V_{i})V_{i}$
is $\epsilon$ close to $\nabla_{\Sigma}f(K,\Sigma)$ with high probability under some conditions. We break $\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)$ into two terms,
$\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)$ $=(\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)-\nabla_{\Sigma}f(K,\Sigma))+(\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma))$ (58)
For the first term, by Lemma $4.8$ we have $\|\nabla_{\Sigma}f(K,\Sigma+V)-\nabla_{\Sigma}f(K,\Sigma)\|<\frac{\epsilon}{2}$ if $r_{2}\leq\frac{\epsilon}{2h_{9}}$. And because $\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)=\mathbb{E}[\nabla_{\Sigma}f(K,\Sigma+V)]$, we have $\|\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)-\nabla_{\Sigma}f(K,\Sigma+V)\|_{F}\leq\frac{\epsilon}{2}$ by triangle inequality. For the second term $\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)$. Consider a single sample $V_{j}$ from the distribution. Assume $f(K,\Sigma+V_{j})\leq 2f(K,\Sigma)$, then $\|\frac{n}{r_{2}^{2}}f(K,\Sigma+V_{j})V_{j}\|\leq\frac{2nf(K,\Sigma)}{r_{2}}$. Also, note that $\mathbb{E}\left[\frac{n}{r_{2}^{2}}f(K,\Sigma+V_{j})V_{j}\right]=\mathbb{E}\left[\widetilde{\nabla}_{\Sigma}\right]=\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)$. Now, given that $\mathbb{E}\left[\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\right]=0$, $\|\frac{n}{r_{2}^{2}}f(K,\Sigma+V_{j})V_{j}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\|_{F}\leq\frac{2nf(K,\Sigma)}{r_{2}}+\frac{\epsilon}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$, and if $\mathbb{E}[\|\frac{n}{r^{2}}f(K,\Sigma+V_{j})V_{j}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\|_{F}^{2}]=\sigma_{V}^{2}\leq\left(\frac{2nf(K,\Sigma)}{r_{2}}\right)^{2}+\left(\frac{\epsilon}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}\right)^{2}$ then by the Vector Bernstein Inequality we have $\mathbb{P}\big(\big\|\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f_{r_{2}}(K,\Sigma)\big\|\geq\frac{\epsilon}{2}\big)\leq n\cdot\exp{\big(-\frac{3}{2}\cdot\frac{\frac{\epsilon^{2}}{4}M}{3\sigma_{V}^{2}+R_{2}\frac{\epsilon}{2}}\big)}\leq\kappa_{2}$, $R_{2}=\frac{2nf(K,\Sigma)}{r_{2}}+\frac{\epsilon}{2}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$, and $\kappa_{2}\in(0,1)$. This also gives us that the minimum samples: $M\geq\log\left(\frac{d}{\kappa_{2}}\right)\cdot\frac{4}{\epsilon^{2}}\left(2\sigma_{V}^{2}+R_{2}\frac{\epsilon}{3}\right)$.
In the above, we have demonstrated that $\widetilde{\nabla}_{\Sigma}$ is $\epsilon$ close to $\nabla_{\Sigma}f(K,\Sigma)$ under some conditions. Now we attempt to prove that $\widehat{\nabla}_{K}$ is $\epsilon$ close to $\nabla_{\Sigma}f(K,\Sigma)$ with high probability $1-\kappa_{2}$, under the conditions given in the lemma. Define
$\nabla_{\Sigma}^{\prime}:=\frac{1}{M}\sum_{i=1}^{M}\frac{n}{r^{2}}f^{(l)}(K,\Sigma+V_{j})V_{j}.$
We break $\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)$ into three terms,
$\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)=(\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}^{\prime})+(\nabla_{\Sigma}^{\prime}-\hat{\nabla}_{\Sigma})+(\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma))$
For the first term, we have
$\sum_{j=0}^{l-1}Q(x_{j}^{i})^{2}+(u_{j}^{i})^{\top}R(u_{j}^{i})+\tau\log\pi(u_{j}|x_{j})$ $\leq\Gamma\mathbb{E}\left[\sum_{j=0}^{l-1}Qx_{j}^{2}+u_{j}^{\top}Ru_{j}+\tau\log\pi(u_{j}|x_{j})\right]$
$\leq\Gamma\mathbb{E}\left[\sum_{j=0}^{\infty}Qx_{j}^{2}+u_{j}^{\top}Ru_{j}+\tau\log\pi(u_{j}|x_{j})\right]$
$\leq\Gamma L^{2}f(K,\Sigma+V_{i})$
$\leq 2\Gamma L^{2}f(K,\Sigma)$
Where $|x_{0}^{i}|\leq L$. Then since $\|\nabla_{\Sigma}^{\prime}\|\leq\frac{\epsilon}{3}+\|\widehat{\nabla}_{\Sigma}\|\leq\frac{5\epsilon}{6}+\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}$ so the norm of each sample is bounded, $\mathbb{E}[\widehat{\nabla}_{\Sigma}-\nabla^{\prime}_{\Sigma}]$ = 0, and assuming the variance is bounded as well then by the Vector Bernstein Inequality $\mathbb{P}\left(\|\widehat{\nabla}_{\Sigma}-\nabla_{\Sigma}^{\prime}\|>\frac{\epsilon}{3}\right)\leq d\cdot\exp{\left(-\frac{3}{2}\cdot\frac{\frac{\epsilon^{2}}{9}M}{3\sigma_{V^{\prime}}^{2}+R_{2}^{\prime}\frac{\epsilon}{3}}\right)}\leq\kappa_{2}$ if $M\geq\frac{6}{\epsilon^{2}}\log\left(\frac{d}{\kappa_{2}}\right)\left(3\sigma_{V^{\prime}}^{2}+R_{2}^{\prime}\frac{\epsilon}{3}\right)$, so $\|\widehat{\nabla}_{\Sigma}-\nabla^{\prime}_{\Sigma}\|\leq\frac{\epsilon}{3}$ with probability at least $1-\kappa_{2}$. For the second term, we have
[…]
$\leq\frac{1}{M}\cdot\frac{n}{r_{2}}\sum_{i=1}^{M}\gamma^{l}\Bigg|\left(|Q|+\|K\|^{2}\|R\|\right)\frac{f(K,\Sigma+V_{i})-\left(Tr((\Sigma+V_{i})R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma+V_{i}|)\right)}{|Q|}\Bigg|$
$+\Bigg|\frac{Tr((\Sigma+V_{i})R)-\frac{\tau}{2}(n+\log(2\pi)^{n}|\Sigma+V_{i}|)}{1-\gamma}\Bigg|$
$\leq\frac{1}{M}\cdot\frac{n}{r_{2}}\sum_{i=1}^{M}\gamma^{l}\left[\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}\right)2f(K,\Sigma)\right$.
$\quad+\left.\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\right)\left|Tr((\Sigma+V_{i})R)-\frac{\tau}{2}\left(n+2\log(2\pi)^{n}|\Sigma|\right)\right|\right]$
$\leq\frac{1}{M}\cdot\frac{n}{r_{2}}M\gamma^{l}\big[\big(1+\frac{\|K\|^{2}\|R\|}{|Q|}\big)2f(K,\Sigma)$
$\quad+\big(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\big)\left(\left|2Tr(\Sigma R)\right|+\frac{\tau}{2}\left(n+2\log(2\pi)^{n}|\Sigma|\right)\right)\big]$
$\leq\frac{\epsilon}{3},$
if $|Tr(\Sigma+V_{i})|\leq 2|Tr(\Sigma)|$, $\|\Sigma+V_{i}\|\leq 2\|\Sigma\|$, and $l\geq$
$\frac{\log\left(\frac{\epsilon r_{2}}{3n}\right){\log(\gamma)}-\log\left[\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}\right)2f(K,\Sigma)+\left(1+\frac{\|K\|^{2}\|R\|}{|Q|}+\frac{1}{1-\gamma}\right)(|\text{Tr}(\Sigma R)|+\frac{\tau}{2}|(n+2\log(2\pi)^{n}|\Sigma|)|)\right]}{\log(\gamma)}$
For the third term, based on the previous proof, we have $\|\widetilde{\nabla}_{\Sigma}-\nabla_{\Sigma}f(K,\Sigma)\|\leq\frac{\epsilon}{3}$ with high probability $1-\kappa_{2}$ under conditions given in the lemma. ∎
C.10 Proof of Lemma 4.9
Proof.
Note that if $r=\min\{\frac{S_{K,\Sigma}}{2h_{K}},h_{\Sigma}\}$, we have $\frac{S_{K,\Sigma}}{2}\leq S_{K+U_{i},\Sigma}\leq\frac{3S_{K,\Sigma}}{2}$. Let $\widetilde{S}=\frac{1}{M}\sum_{i=1}^{M}S_{K+U_{i},\Sigma}$ and $\widetilde{S}^{(l)}=\frac{1}{M}\sum_{i=1}^{M}S^{(l)}_{K+U_{i},\Sigma}$ We broke $\widehat{S}-S_{K,\Sigma}$ into the following three terms:
$\widehat{S}-{S_{K,\Sigma}}=\widehat{S}-\widetilde{S}^{(l)}+\widetilde{S}^{(l)}-\widetilde{S}+\widetilde{S}-\ {S_{K,\Sigma}}$
[…]
For the bound on $\widehat{S}$, apply Weyl’s Theorem when $\epsilon\leq\mu/2$ we have $\sigma_{min}(\widehat{S})\geq\sigma_{min}(S_{K,\Sigma})-\mu/2\geq\mu/2$
C.11 Proof of Lemma 4.10
Proof.
By triangle inequality we have:
$|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma}|\leq|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma^{\prime}}|+|q_{K,\Sigma^{\prime}}-q_{K,\Sigma}|$
For the first term, when $\|K^{\prime}-K\|\leq\min\{h_{\Sigma},\|K\|\}$ and $\|\Sigma^{\prime}-\Sigma\|\leq\|\Sigma\|$, using lemma 4.4 we have
$|q_{K^{\prime},\Sigma^{\prime}}-q_{K,\Sigma^{\prime}}|$ $=\frac{\text{Tr}(\Sigma^{\prime}\gamma(P_{K}-P_{K}^{\prime})(B^{\top}B+D^{\top}D))}{1-\gamma}$
$\leq\frac{\gamma\|\Sigma^{\prime}\|\|B^{\top}B+D^{\top}D\|}{1-\gamma}|P_{K}-P_{K^{\prime}}|$
$\leq\frac{2\gamma\|\Sigma\|\|B^{\top}B+D^{\top}D\|}{1-\gamma}h_{5}\|K^{\prime}-K\|$
$:=h_{10}\|K^{\prime}-K\|$
For the second term, from intermediate step of lemma 3.3, we have:
$|q_{K,\Sigma^{\prime}}-q_{K,\Sigma}|$ $\leq\frac{m}{2}Tr((\Sigma^{-1}\Sigma^{\prime}-I)^{2})+Tr\left(\nabla_{\Sigma}q_{K,\Sigma}^{\top}(\Sigma-\Sigma^{\prime})\right)$
$\leq\frac{m}{2}\|\Sigma^{-1}\Sigma^{\prime}-I\|_{F}^{2}+\|\nabla_{\Sigma}q_{K,\Sigma}\|\|\Sigma^{\prime}-\Sigma\|$
$\leq\frac{m\|\Sigma^{-1}\|_{F}}{2}\|\Sigma^{\prime}-\Sigma\|_{F}^{2}+\overline{\|\nabla_{\Sigma}q_{K,\Sigma}\|}\|\Sigma^{\prime}-\Sigma\|_{F}$
$:=h_{11}\|\Sigma^{\prime}-\Sigma\|_{F}$
Combining the above completes the proof. ∎
Appendix D Standard Matrix Perturbation and Concentrations
In this section, we review several basic matrix tools that were used throughout the paper.
D.1 Vector Bernstein inequality
Lemma D.1.
Let $\{Z_{i}\}_{i=1}^{N}$ be a set of $N$ independent random vectors of dimension n{n} with $\mathbb{E}[Z_{i}]=Z$}, $\|Z_{i}-Z\|\leq R_{Z}$} almost surely, and maximum variance $\left\|\mathbb{E}(Z_{i}^{\intercal}Z_{i})-Z^{\intercal}Z\right\|\leq\sigma_{Z}^{2}$}, and sample average $\widehat{Z}:=\frac{1}{N}\sum_{i=1}^{N}Z_{i}$}. Let a small tolerance $\epsilon\geq 0$} and small probability $0\leq\kappa\leq 1$} be given. If
$N\geq\frac{2n}{\epsilon^{2}}\left(\sigma_{Z}^{2}+\frac{R_{Z}\epsilon}{3\sqrt{n}}\right)\log\left[\frac{n}{\mu}\right]$
then
$\mathbb{P}\left[\left\|\widehat{Z}-Z\right\|_{F}\leq\epsilon\right]\geq 1-\mu.$ (59)
Proof.
This Lemma is directly obtained by applying Lemma C.6 in [10] to the case of vectors. ∎
D.2 Weyl’s Inequality for singular values
Suppose $B=A+E$, then the singular values of $B$ are within E to the corresponding singular values of A. In particular, $\|B\|\leq\|A\|+\|E\|$ and $\sigma_{min}(B)\geq\sigma_{min}(A)-\|E\|$ .
D.3 Perturbation of Inverse
Let $B=A+E$, suppose $E\leq\sigma_{min}(A)/2$, then $\|B^{-1}-A^{-1}\|\leq 2\|A-B\|/\sigma_{min}(A)$
D.4 Matrix Norm
For matrix $A,B\in\mathbb{R}^{n\times m}$, we have $\|A^{-1}\|\geq\|A\|^{-1}$ and $|Tr(A^{\top}B)|\leq\|A^{\top}\||Tr(B)|=\|A\||Tr(B)|$. If $A\succeq 0$, we have $Tr(A)\geq\|A\|$